### v1

# SABR-Informed Multitask Gaussian Process: A Synthetic-to-Real Framework for Implied Volatility Surface Construction

Abstract
Constructing the Implied Volatility Surface (IVS) is a challenging task in quantitative finance due to the complexity of real markets and the sparsity of market data. Structural models like Stochastic Alpha Beta Rho (SABR) model offer interpretability and theoretical consistency but lack flexibility, while purely data-driven methods such as Gaussian Process regression can struggle with sparse data. We introduce SABR-Informed Multi-Task Gaussian Process (SABR-MTGP), treating IVS construction as a multi-task learning problem. Our method uses a dense synthetic dataset from a calibrated SABR model as a source task to inform the construction based on sparse market data (the target task). The MTGP framework captures task correlation and transfers structural information adaptively, improving predictions particularly in data-scarce regions. Experiments using Heston-generated ground truth data under various market conditions show that SABR-MTGP outperforms both standard Gaussian process regression and SABR across different maturities. Furthermore, an application to real SPX market data demonstrates the method’s practical applicability and its ability to produce stable and realistic surfaces. This confirms our method balances structural guidance from SABR with the flexibility needed for market data.
keywords:
Gaussian Process , Implied Volatility , Multi-Task Learning , SABR , Machine Learning.
[…]
[inst1]organization=Department of Mathematics, University of Macau,addressline=Avenida da Universidade Taipa, postcode=999078, state=Macau, country=China
1 Introduction
Constructing the Implied Volatility Surface (IVS) is an important and difficult task in quantitative finance. A common difficulty arises from sparse market option data, especially for long-dated expiration or strikes far from the money. This sparsity causes difficulties for conventional construction approaches.
Structural models, such as the SABR model [1], are popular because they capture volatility smiles and skews with few parameters. The SABR model assume that the underlying asset and its volatility follow specific stochastic differential equations. While these models provide reasonable interpolations, their predetermined functional forms may suffer from a lack of flexibility and not fully reflect market complexities. In contrast, data-driven approaches offer more flexibility, adapting to market patterns without imposing rigid structures. Gaussian Processes (GP) [2], provide a Bayesian framework for regression. However, these methods typically require substantial data to achieve reliable results. With sparse observations, GP may fail to capture patterns of implied volatility.
In order to address the challenges in data-driven approaches, recent works discuss how to incorporate machine learning methods with prior financial knowledge. Cousin et al. [3], Chataigner et al. [4], Ackerer et al. [5], Zheng et al. [6], Gonon et al. [7], Hoshisashi et al. [8] present methods that imposing constraints derived from financial theory within a neural network or a Gaussian Process. In contrast, Chen et al. [9] use typical transfer learning methods for option pricing. A neural network is trained on a structural model’s synthetic data (generated by the Black-Scholes model) is primarily, and it provide a good initialization for subsequent fine-tuning on empirical data.
Different to their methods, we present the SABR-Informed Multi-Task Gaussian Process (SABR-MTGP) which takes advantage of both structural and data-driven models. The key idea is to treat the SABR-generated IVS as an information source (a ‘source’ task) and market observations as the primary target task. Then, a multi-task learning framework [10, 11] is used where the relationship between the theoretical structure (SABR) and empirical observations (market data) is learned adaptively during a joint optimization process. The implementation consists of two main stages. First, we generate a synthetic dataset using the calibrated SABR model. This dataset embodies smile and term-structure characteristics typical of SABR dynamics. Second, we train a multi-task Gaussian Process (MTGP) model simultaneously on these dense synthetic data and the sparse market observations. The MTGP framework learns correlation patterns between tasks through shared and task-specific covariance components. We use task embeddings with hierarchical regularization to appropriately balance how much structural guidance from SABR is used when constructing the representation from real market data. This avoids both overly rigid adherence to SABR patterns and complete disregard for its structural insights.
We evaluated our method through numerical experiments using the Heston model [12] to generate sparse ‘market’ datasets. Comparing SABR-MTGP against standard GP and SABR with interpolation, we find that our approach generally yields more accurate volatility surfaces across diverse market conditions and maturities, with particular advantages in sparse data regions. We further demonstrate the practical applicability of our method through a case study on real SPX market data. These results show the effectiveness of our method in integrating financial domain knowledge within a flexible machine learning framework.
Related Literature
Gaussian processes have found increasing application across finance and economics due to their flexibility and ability to quantify uncertainty. A significant body of work focuses on derivative pricing, valuation, and hedging. GP models have been used for constructing implied volatility curves [13, 14], constructing financial term structures [15], accelerating pricing and hedging calculations for various options [13, 16, 17, 18, 19], modeling derivative portfolios for CVA computation [20], pricing complex insurance products like variable annuities [21], and approximating Greeks for hedging [22]. Beyond direct pricing, GP models serve as flexible tools for calibrating implied and local volatility surfaces [14, 4] or measuring portfolio tail risk [23].
Furthermore, GP models are employed in broader financial and economic time series analysis, including imputing missing financial data [24], forecasting real estate prices [25], modeling inflation dynamics [26], understanding determinants of carbon market prices [27], and developing nonparametric vector autoregressions for macroeconomic analysis [28].
While these studies highlight the versatility of GP models, our work specifically contributes by employing a multi-task learning framework informed by a structural model (SABR) to address the challenge of sparse data in IVS construction, learning the relationship between the structural prior and market data adaptively.
The remainder of the paper is organized as follows: Section 2 introduces some basic concept in implied volatility surface and Gaussian Process. Section 3 provides the details of the proposed methods. In Section 4, we describe the setup of the numerical experiments and the results are presented in Section 5. In Section 6, we demonstrate the performance of the proposed method in real market data.
2 Background
This section provides a necessary background on the implied volatility surface and the Gaussian process.
2.1 Implied Volatility Surface
The Black-Scholes model [29] is a basic framework for European option pricing under specific assumptions: key assumptons include that the underlying asset price follows the Geometric Brownian Motion, and asset volatility are constants. Observed market option prices frequently deviate from the prices given by the Black-Scholes model under the assumption of constant volatility. This difference is captured through the Implied Volatility (IV), a way to match the theoretical framework with market observations. For an observed market price of call options $C_{mkt}(K,\tau)$ with time-to-maturity $\tau$ and strike price $K$, the corresponding IV $\sigma_{mkt}(K,\tau)$ is the unique positive value that makes the following equation holds:
$C_{mkt}(K,\tau)=C_{BS}(K,\tau,\sigma=\sigma_{mkt}(K,\tau)\,;\theta_{BS}),$ (1)
where $C_{BS}$ is the Black-Scholes formula, and $\theta_{BS}$ is the collection of rest parameters in the Black-Scholes formula. Finding $\sigma_{mkt}(K,\tau)$ needs solving the Black-Scholes formula backwards, usually done with iterative root-finding algorithms (like Newton-Raphson). When $\sigma_{mkt}(K,\tau)$ is shown as a function of both variables, the resulting three-dimensional surface is the Implied Volatility Surface (IVS). The patterns of IVS reflect market participants’ risk preferences and expectations about future price dynamics beyond the standard Black-Scholes assumptions. Accurate IVS construction is important for derivative pricing, hedging strategies, and risk management across option portfolios.
2.2 Gaussian Process
A Gaussian Process (GP) is a Bayesian non-parametric approach that can be used in regression problems [2]. A GP defines a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$ such that any finite set of function values has a joint Gaussian distribution. A GP is fully specified by its mean function $m:\mathbb{R}^{d}\to\mathbb{R}$ and covariance function (or kernel) $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$:
$f\sim\mathcal{GP}(m,k).$
The mean function is often set to zero for simplicity. The kernel encodes assumptions about function properties like smoothness or characteristic length-scales, defining the covariance between function values at any two input points. A commonly used kernel is the Matérn 5/2 kernel function which is a second-order differentiable kernel function:
[…]
$\text{for}\quad\mathbf{x,x^{\prime}}\in\mathbb{R}^{d}$, where $\sigma_{f}$ is the scaling hyperparameter and $\gamma$ is the length-scales hyperparameter.
Given training data $\mathcal{D}=\{(\mathbf{X,y})\}$ where the components in $\mathbf{y}=[y_{1},\dots,y_{N}]^{\intercal}$ are noisy observations at $N$ locations $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{N}]^{\intercal}$. $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ where $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{n}^{2})$ is noise. As a result, a predictive distribution for the function value $f_{\ast}$ at a new test point $\mathbf{x}_{\ast}$ is given by the posterior $p\big{(}f_{\ast}(\mathbf{x}_{\ast})|\mathcal{D}\big{)}\sim\mathcal{N}(\mu_{\ast},\sigma_{\ast})$, where
$\mu_{\ast}$ $=k(\mathbf{X,x_{\ast}})^{\intercal}\big{(}k(\mathbf{X,X})+\sigma_{n}^{2}\mathbf{I}\big{)}^{-1}\mathbf{y},$ (2)
$\sigma_{\ast}^{2}$ $=k(\mathbf{x}_{\ast},\mathbf{x}_{\ast})-k(\mathbf{X,x_{\ast}})^{\intercal}\big{(}k(\mathbf{X,X})+\sigma_{n}^{2}\mathbf{I}\big{)}^{-1}k(\mathbf{X,x_{\ast}}).$ (3)
Here, $k(\mathbf{X,x_{\ast}})\coloneqq\left[k(\mathbf{x}_{1},\mathbf{x_{\ast}}),\dots,k(\mathbf{x}_{N},\mathbf{x_{\ast}})\right]^{\intercal}$, $\mathbf{I}$ is an identity matrix and
$k(\mathbf{X,X})\coloneqq\begin{bmatrix}k(\mathbf{x}_{1},\mathbf{x}_{1})&\dots&k(\mathbf{x}_{1},\mathbf{x}_{N})\\ \vdots&\ddots&\vdots\\ k(\mathbf{x}_{N},\mathbf{x}_{1})&\dots&k(\mathbf{x}_{N},\mathbf{x}_{N})\end{bmatrix}.$
The kernel hyperparameters and noise variance $\sigma_{n}^{2}$ are usually estimated by maximizing the marginal likelihood of the data. This provides an automatic way to balance data fit and model complexity [30]. GP provide a Bayesian framework that can model complex, non-linear relationships without imposing rigid functional forms. However, they can face challenges when training data is sparse, as often happens with implied volatility observations.
3 SABR-Informed Multi-Task Gaussian Process
In this section, we describe our approach for constructing the implied volatility surface by combining SABR structure with GP flexibility.
3.1 Problem Formulation
The primary objective of this work is to construct an accurate Implied Volatility Surface (IVS). The IVS, denoted as $\sigma(K,\tau)$, is a function of the option strike price $K$ and its time-to-maturity $\tau$. We can represent the input features as a two-dimensional vector $\mathbf{x}=[K,\tau]^{\intercal}$. The task is thus to learn a regression model $f:\mathbb{R}^{2}\to\mathbb{R}^{+}$ that maps these input features to the corresponding implied volatility, i.e., $y=f(\mathbf{x})=\sigma(K,\tau)$.
Suppose that we have $N$ observed locations $\mathbf{X}_{\mathcal{T}}=[\mathbf{x}_{\mathcal{T},1},\dots,\mathbf{x}_{\mathcal{T},N}]^{\intercal}$, $\mathbf{x}_{\mathcal{T},i}=[K_{i},\tau_{i}]^{\intercal}$, and $\mathbf{y}_{\mathcal{T}}=[y_{\mathcal{T},1},\dots,y_{\mathcal{T},N}]^{\intercal}$ contains the corresponding observed market implied volatilities, $y_{\mathcal{T},i}=\sigma_{\rm mkt}(K_{i},\tau_{i})$. We denote the market-observed option dataset as $\mathcal{D}_{\mathcal{T}}=\{(\mathbf{X}_{\mathcal{T}},\mathbf{y}_{\mathcal{T}})\}$. A significant challenge in practice is that this market data $\mathcal{D}_{\mathcal{T}}$ is often sparse, particularly for options with long maturities or strikes far from the current underlying price (deep out-of-the-money or in-the-money). This sparsity can make it difficult for purely data-driven models to construct reliable IVS.
To address this challenge, our methodology transfers information from a structural financial model, specifically the SABR model. The core idea is to generate a dense synthetic dataset, denoted $\mathcal{D}_{\mathcal{S}}$, using a calibrated SABR model. This synthetic dataset acts as a source of information to guide the learning process and prediction, especially in regions where market data is scarce. The subsequent sections will show how this synthetic data is generated and incorporated within a multi-task learning framework.
3.2 Synthetic Data Generation via SABR
A key step in our work was creating a synthetic dataset that captures structural patterns of volatility surfaces. In this paper, we used the SABR model [1] for incorporating financial theory to guide the model. SABR is one of the most widely used parametric models in quantitative finance for modeling implied volatility smiles and skews. Hagan’s asymptotic expansion provides a closed-form approximation for implied volatility under SABR, which is crucial for efficiently generating the dense synthetic dataset $\mathcal{D}_{\mathcal{S}}$.
The SABR model describes the dynamics of the forward price $F_{t}\coloneqq S_{t}e^{(r-q)(T-t)}$ (where $S_{0}$ is the current price of underlying asset, $r$ is the risk-free interest rate, $q$ is the dividend rate, $T$ is the maturity) and its instantaneous volatility $\alpha_{t}$ via stochastic differential equations:
$dF_{t}$ $=\alpha_{t}F_{t}^{\beta}dW_{t}^{(1)},$ (4)
$d\alpha_{t}$ $=\nu\alpha_{t}dW_{t}^{(2)},$ (5)
$dW_{t}^{(1)}dW_{t}^{(2)}$ $=\rho dt.$ (6)
Here, $\beta\in[0,1]$ relates volatility to the price level, $\nu$ is the volatility-of-volatility, and $\rho$ is the correlation between two Brownian motions. Each parameter influences the IVS shape: $\alpha$ determines the volatility level, $\beta$ affects the backbone shape, $\rho$ controls the skew, and $\nu$ drives smile convexity.
SABR Model Asymptotic Expansion
Practical applications often use Hagan’s asymptotic expansion [1], which provides closed-form approximations for implied volatility $\sigma_{\rm SABR}(K,\tau,F;\alpha,\beta,\rho,\nu)$.
$\sigma_{\rm SABR}(K,\tau,F)=\left\{\begin{aligned} &I^{0}(K,\tau,F)\left[1+I^{1}(K,F)\tau\right]\quad&\text{if}\,K\neq F,\\ &\frac{\alpha}{F^{1-\beta}}\left[1+\left(\frac{(1-\beta)^{2}\alpha^{2}}{24F^{2-2\beta}}+\frac{\rho\beta\alpha\nu}{4F^{1-\beta}}+\frac{2-3\rho^{2}}{24}\nu^{2}\right)\tau\right]\quad&\text{if}\,K=F.\end{aligned}\right.$ (7)
where:
$I^{0}(K,\tau,F)=$ $\left(\frac{\alpha z(K,F)}{\chi(z)(FK)^{(1-\beta)/2}}\right)\left[1+\frac{(1-\beta)^{2}}{24}\log^{2}(F/K)+\frac{(1-\beta)^{4}}{1920}\log^{4}(F/K)\right]^{-1},$
$I^{1}(K,F)=$ $\frac{(1-\beta)^{2}\alpha^{2}}{24(FK)^{1-\beta}}+\frac{\rho\beta\nu\alpha}{4(FK)^{(1-\beta)/2}}+\frac{2-3\rho^{2}}{24}\nu^{2},$
$z(K,F)=$ $\frac{\nu}{\alpha}(FK)^{(1-\beta)/2}\log(F/K),\quad\chi(z)=\log\left[\frac{\sqrt{1-2\rho z+z^{2}}+z-\rho}{1-\rho}\right].$
Let $\mathbb{T}_{\text{mkt}}=\{\tau_{(1)},\dots,\tau_{(P)}\}$ be the set of unique maturities in the market data $\mathcal{D}_{\mathcal{T}}$, ordered $\tau_{(1)}<\tau_{(2)}<\dots<\tau_{(P)}$. For each maturity slice $\tau_{(p)}$, we calibrated SABR parameters $(\alpha_{(p)},\rho_{(p)},\nu_{(p)})$ to match the observed market smile by minimizing the squared errors by Nelder-Mead algorithm [31]:
$\min_{\alpha_{(p)},\rho_{(p)},\nu_{(p)}}\sum_{i:\,\tau_{i}=\tau_{(p)}}\left(\sigma_{\rm SABR}(K_{i},F_{\tau_{(p)}},\tau_{(p)};\,\alpha_{(p)},\beta,\rho_{(p)},\nu_{(p)})-y_{\mathcal{T},i}\right)^{2}$ (8)
Here, $F_{\tau_{(p)}}=S_{0}e^{(r-q)\tau_{(p)}}$ is the forward price at maturity $\tau_{(p)}$. We fixed $\beta=0.5$ throughout our experiments, a common choice for equity options. Calibrating all four SABR parameters ($\alpha,\beta,\rho,\nu$) simultaneously for each maturity slice can be challenging and numerically unstable, especially with sparse or noisy data. Calibrating $\beta$ often requires richer data across strikes than might be available, particularly for longer maturities. Fixing $\beta$ reduces the dimensionality of the optimization problem for each slice from four parameters to three ($\alpha,\rho,\nu$), significantly improving the stability and speed of the calibration process. The parameters $\alpha_{(p)}$, $\rho_{(p)}$, and $\nu_{(p)}$ were optimized within realistic ranges: $2.00.01<\alpha_{(p)}<$ 2.0, $-0.99<\rho_{(p)}<0.0$, and $1.50.05<\nu_{(p)}<$ 1.5. This yielded calibrated parameters $\{(\alpha_{(p)},\rho_{(p)},\nu_{(p)})\}_{p=1}^{P}$ for each observed maturity slice.
To generate the synthetic dataset, we needed SABR parameters for any desired maturity. We defined a dense grid of $M$ points $\mathbf{X}_{\mathcal{S}}=[\mathbf{x}_{\mathcal{S},1},\dots,\mathbf{x}_{\mathcal{S},M}]^{\intercal}$, where $\mathbf{x}_{\mathcal{S},j}=[\tilde{K}_{j},\tilde{\tau}_{j}]^{\intercal}$, covering the region of interest.
For any grid maturity $\tilde{\tau}$, we determined the SABR parameters as follows:
  • If $\tilde{\tau}$ matches a calibrated maturity $\tau_{(p)}$, we used the calibrated parameters $(\alpha_{(p)},\rho_{(p)},\nu_{(p)})$.
  • If $\tilde{\tau}$ is between two calibrated maturities, $\tau_{(p)}<\tilde{\tau}<\tau_{(p+1)}$, we used piecewise linear interpolation to get parameters $(\alpha_{\tilde{\tau}},\rho_{\tilde{\tau}},\nu_{\tilde{\tau}})$. For example, for $\alpha$:
    $\alpha_{\tilde{\tau}}=\alpha_{(p)}+\frac{\tilde{\tau}-\tau_{(p)}}{\tau_{(p+1)}-\tau_{(p)}}(\alpha_{(p+1)}-\alpha_{(p)})$ (9)
    The same applied for $\rho_{\tilde{\tau}}$ and $\nu_{\tilde{\tau}}$.
  • For extrapolation beyond the calibrated range ($\tilde{\tau}<\tau_{(1)}$ or $\tilde{\tau}>\tau_{(P)}$), we used constant extrapolation (parameters from the nearest calibrated maturity).
With parameters for each grid point, we computed synthetic volatilities
$y_{\mathcal{S},j}=\sigma_{\rm SABR}(\tilde{K}_{j},F_{\tilde{\tau}_{j}},\tilde{\tau}_{j};\alpha_{\tilde{\tau}_{j}},\beta,\rho_{{\tilde{\tau}}_{j}},\nu_{{\tilde{\tau}}_{j}}).$
This produced the synthetic source dataset $\mathcal{D}_{\mathcal{S}}=\{(\mathbf{X}_{\mathcal{S}},\mathbf{y}_{\mathcal{S}})\}$, where $\mathbf{y}_{\mathcal{S}}=[y_{\mathcal{S},1},\dots,y_{\mathcal{S},M}]^{\intercal}$. In practice, Gaussian noise $\epsilon_{\mathcal{S}}\sim\mathcal{N}(0,\sigma^{2}_{\text{syn}})$ is added to the synthetic volatilities to prevent the model from becoming overly dependent on the synthetic data.
3.3 Multi-Task Gaussian Process Formulation
To combine information from the dense synthetic data (SABR-generated) and sparse market observations, we used a Multi-Task Gaussian Process (MTGP) framework. Standard GP models model a single output, while MTGP models handle multiple related outputs by capturing correlations between tasks [10, 11].
We designated the synthetic SABR data $\mathcal{D}_{\mathcal{S}}$ as the source task and the sparse market data $\mathcal{D}_{\mathcal{T}}$ as the target task. For each task, observed volatilities were modeled as noisy realizations of a latent function:
$y_{\mathcal{S}}(\mathbf{x})=f_{\mathcal{S}}(\mathbf{x})+\varepsilon_{\mathcal{S}},\quad\varepsilon_{\mathcal{S}}\sim\mathcal{N}(0,\sigma^{2}_{\mathcal{S}}),\quad y_{\mathcal{T}}(\mathbf{x})=f_{\mathcal{T}}(\mathbf{x})+\varepsilon_{\mathcal{T}},\quad\varepsilon_{\mathcal{T}}\sim\mathcal{N}(0,\sigma^{2}_{\mathcal{T}})$ (10)
The noise variances $\sigma^{2}_{\mathcal{S}}$ and $\sigma^{2}_{\mathcal{T}}$ reflect different uncertainty levels in the synthetic SABR approximations and market observations.
A key idea was decomposing each latent function to facilitate information sharing. We expressed each task’s function as:
$f_{\mathcal{S}}(\mathbf{x})=g_{\mathcal{S}}(\mathbf{x})+h_{\mathcal{S}}(\mathbf{x}),\quad f_{\mathcal{T}}(\mathbf{x})=g_{\mathcal{T}}(\mathbf{x})+h_{\mathcal{T}}(\mathbf{x})$ (11)
This decomposition included:
  • Task-specific components $g_{\mathcal{S}}$ and $g_{\mathcal{T}}$ that captured unique patterns in each task
  • Shared components $h_{\mathcal{S}}$ and $h_{\mathcal{T}}$ that contributed to the construction of correlations between tasks
We placed independent GP priors on the task-specific components, using a common input covariance function $k$ (e.g., a Matérn 5/2 kernel) with task-specific variance scaling parameters:
$g_{\mathcal{S}}\sim\mathcal{GP}(0,\kappa_{\mathcal{S}}^{2}k),\quad g_{\mathcal{T}}\sim\mathcal{GP}(0,\kappa_{\mathcal{T}}^{2}k)$ (12)
The parameters $\kappa_{\mathcal{S}}^{2}$ and $\kappa_{\mathcal{T}}^{2}$ control the amount of task-specific variance.
For the shared components, we use learned task embeddings to infer the relationship. We modeled $h_{\mathcal{S}}$ and $h_{\mathcal{T}}$ as jointly Gaussian with a covariance structure:
$\text{Cov}(h_{\mathcal{Z}}(\mathbf{x}),h_{\mathcal{Z}^{\prime}}(\mathbf{x}^{\prime}))=\underbrace{\sigma^{2}_{h}\exp\left(-\frac{\|\mathbf{e}_{\mathcal{Z}}-\mathbf{e}_{\mathcal{Z}^{\prime}}\|^{2}}{l_{h}^{2}}\right)}_{\eqqcolon\,\hat{C}_{\mathcal{Z},\mathcal{Z}^{\prime}}}k(\mathbf{x},\mathbf{x}^{\prime})\quad\text{for}\quad\mathcal{Z,Z^{\prime}}\in\{\mathcal{S,T}\}$ (13)
The term $\hat{C}_{\mathcal{Z},\mathcal{Z}^{\prime}}$ depends on task embeddings $\mathbf{e}_{\mathcal{Z}},\mathbf{e}_{\mathcal{Z}^{\prime}}\in\mathbb{R}^{d^{\prime}}$ ($d^{\prime}$ is a hyperparemeter that need to be predetermined) , a shared variance scaling parameter $\sigma^{2}_{h}$ and a length-scale parameter $l_{h}$. This allowed the model to learn the level of information sharing. Tasks with closer embeddings would have stronger correlation through the shared component. With components $g_{\mathcal{S}},g_{\mathcal{T}},h_{\mathcal{S}},h_{\mathcal{T}}$, the overall covariance structure follows an Intrinsic Coregionalization Model (ICM) [10]:
$\text{Cov}(f_{\mathcal{Z}}(\mathbf{x}),f_{\mathcal{Z}^{\prime}}(\mathbf{x}^{\prime}))=\underbrace{\left(\hat{C}_{\mathcal{Z},\mathcal{Z}^{\prime}}+\kappa^{2}_{\mathcal{Z}}\delta_{\mathcal{Z}\mathcal{Z}^{\prime}}\right)}_{\eqqcolon\,C_{\mathcal{Z},\mathcal{Z}^{\prime}}}k(\mathbf{x},\mathbf{x}^{\prime})\quad\text{for}\quad\mathcal{Z,Z^{\prime}}\in\{\mathcal{S,T}\}$ (14)
Here, $\delta_{\mathcal{Z}\mathcal{Z}^{\prime}}$ is the Kronecker delta, ensuring task-specific components only contribute to their own task’s variance. As a result, the joint distribution of $\mathbf{f}_{\mathcal{S}}=[f_{\mathcal{S}}(\mathbf{x}_{\mathcal{S},1}),\dots,f_{\mathcal{S}}(\mathbf{x}_{\mathcal{S},M})]^{\intercal}$ and $\mathbf{f}_{\mathcal{T}}=[f_{\mathcal{T}}(\mathbf{x}_{\mathcal{T},1}),\dots,f_{\mathcal{T}}(\mathbf{x}_{\mathcal{T},N})]^{\intercal}$ is multivariate Gaussian:
$\begin{bmatrix}\mathbf{f}_{\mathcal{S}}\\ \mathbf{f}_{\mathcal{T}}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mathbf{0}_{M}\\ \mathbf{0}_{N}\end{bmatrix},\mathbf{K}\circ\mathbf{C}\right)$ (15)
where $\circ$ is the Hadamard product (element-wise product). $\mathbf{K}$ and $\mathbf{C}$ are:
$\mathbf{K}=\begin{bmatrix}k(\mathbf{X}_{\mathcal{S}},\mathbf{X}_{\mathcal{S}})&k(\mathbf{X}_{\mathcal{S}},\mathbf{X}_{\mathcal{T}})\\ k(\mathbf{X}_{\mathcal{T}},\mathbf{X}_{\mathcal{S}})&k(\mathbf{X}_{\mathcal{T}},\mathbf{X}_{\mathcal{T}})\end{bmatrix},\quad\mathbf{C}=\begin{bmatrix}C_{\mathcal{S,S}}\mathbf{1}_{M\times M}&C_{\mathcal{S,T}}\mathbf{1}_{M\times N}\\ C_{\mathcal{T,S}}\mathbf{1}_{N\times M}&C_{\mathcal{T,T}}\mathbf{1}_{N\times N}\end{bmatrix}$ (16)
Where $k(\mathbf{X}_{\mathcal{Z}},\mathbf{X}_{\mathcal{Z^{\prime}}})$ is the matrix of kernel evaluations between input points, and $\mathbf{1}_{R\times H}$ is an $R\times H$ matrix of ones.
3.4 Learning Task Relationships via Hierarchical Regularization
The task embeddings $\mathbf{e}_{\mathcal{S}}$ and $\mathbf{e}_{\mathcal{T}}$ are critical. Their locations in the embedding space determine how much structural information transfers between the SABR synthetic data and market observations. However, optimizing these embeddings can be challenging when directly maximizing the marginal likelihood, especially with imbalanced datasets.
When the target dataset $\mathcal{D}_{\mathcal{T}}$ is much sparser than the source dataset $\mathcal{D}_{\mathcal{S}}$ (our typical scenario), standard maximum marginal likelihood estimation (MLE) can lead to two undesirable outcomes:
  1. 1.
    The embeddings might diverge too much, effectively isolating the tasks and preventing useful knowledge transfer from the denser source task.
  2. 2.
    Alternatively, the optimization might be dominated by the dense source data, forcing the embeddings too close together and leading to inappropriate information transfer that ignores potentially significant differences between SABR predictions and market data.
Simply fixing a predetermined correlation level would be too restrictive. We need an adaptive approach that provides guidance while letting the data inform the level of information sharing.
We addressed this using a hierarchical Bayesian formulation within a Maximum A Posteriori (MAP) estimation framework. Instead of treating task embeddings as independent parameters to be learned solely through the likelihood, we assumed they are drawn from a common prior distribution, governed by hyperparameters $\bm{\mu}_{\text{e}}$ and $\sigma^{2}_{\text{e}}$:
$P(\mathbf{e}_{\mathcal{Z}}|\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}})=\mathcal{N}(\mathbf{e}_{\mathcal{Z}}|\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}}\mathbf{I}_{d})\quad\text{for }\mathcal{Z}\in\{\mathcal{S},\mathcal{T}\}.$ (17)
Here, $\bm{\mu}_{\text{e}}\in\mathbb{R}^{d^{\prime}}$ and $\sigma^{2}_{\text{e}}$ controlled the squared distance between task embeddings and their mean. We learned both the embeddings $\mathbf{e}_{\mathcal{S}},\mathbf{e}_{\mathcal{T}}$ and the hyperparameters $\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}}$ from data.
This hierarchical prior was incorporated into the objective function. We aimed to maximize the posterior probability $P(\bm{\theta}|\mathbf{y},\mathbf{X})$, where $\mathbf{y}=[\mathbf{y}_{\mathcal{S}}^{\intercal},\mathbf{y}_{\mathcal{T}}^{\intercal}]^{\intercal}$ and $\bm{\theta}$ represented all model parameters (embeddings, kernel parameters, variances, hierarchical hyperparameters, etc.). By Bayes’ theorem, $P(\bm{\theta}|\mathbf{y},\mathbf{X})\propto P(\mathbf{y}|\mathbf{X},\bm{\theta})P(\bm{\theta})$. Maximizing the posterior is equivalent to minimizing its negative logarithm:
$\hat{\bm{\theta}}_{\text{MAP}}=\arg\min_{\bm{\theta}}\left[-\log P(\mathbf{y}|\mathbf{X},\bm{\theta})-\log P(\bm{\theta})\right]$ (18)
For the prior term $P(\bm{\theta})$, we focused on the hierarchical prior for the embeddings $P(\mathbf{e}_{\mathcal{S}},\mathbf{e}_{\mathcal{T}}|\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}})$. We assumed that other parameters in $\bm{\theta}$ have flat (uninformative) priors and do not affect the optimization minimum. Thus, the relevant negative log prior term is:
$-\log P(\bm{\theta})=$ $-\log P(\mathbf{e}_{\mathcal{S}},\mathbf{e}_{\mathcal{T}}|\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}})$
$=$ $\sum_{\mathcal{Z}\in\{\mathcal{S},\mathcal{T}\}}\left(\frac{\|\mathbf{e}_{\mathcal{Z}}-\bm{\mu}_{\text{e}}\|^{2}}{2\sigma^{2}_{\text{e}}}+\frac{d^{\prime}}{2}\log\sigma^{2}_{\text{e}}+\frac{d^{\prime}}{2}\log(2\pi)\right)$ (19)
The negative log marginal likelihood $-\log P(\mathbf{y}|\mathbf{X},\bm{\theta})$, derived from the joint Gaussian distribution of observations $\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})$ under the MTGP model, is given exactly by:
$-\log P(\mathbf{y}|\mathbf{X},\bm{\theta})=\frac{1}{2}\mathbf{y}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{y}+\frac{1}{2}\log|\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma}|+\frac{M+N}{2}\log(2\pi)$ (20)
where $\mathbf{K}\circ\mathbf{C}$ is the prior covariance matrix of the latent functions at training points ($M$ source, $N$ target), and $\bm{\Sigma}=\text{diag}(\sigma_{\mathcal{S}}^{2}\mathbf{1}_{M},\sigma_{\mathcal{T}}^{2}\mathbf{1}_{N})$ is the diagonal noise covariance matrix.
Substituting the negative log marginal likelihood (NLML, Equation 20) and the relevant negative log prior (Equation 3.4) into the MAP objective (Equation 18), we arrived at the final objective function $\mathcal{L}(\bm{\theta})$ to minimize:
$\mathcal{L}(\bm{\theta})=$ $\underbrace{\frac{1}{2}\mathbf{y}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{y}+\frac{1}{2}\log|\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma}|}_{\text{Data Fidelity Term (NLML)}}$
$\underbrace{+\sum_{\mathcal{Z}\in\{\mathcal{S},\mathcal{T}\}}\left(\frac{\|\mathbf{e}_{\mathcal{Z}}-\bm{\mu}_{\text{e}}\|^{2}}{2\sigma^{2}_{\text{e}}}+\frac{d^{\prime}}{2}\log\sigma^{2}_{\text{e}}\right)}_{\text{Hierarchical Regularization Term (Neg. Log Prior)}}\underbrace{+\frac{M+N+2d^{\prime}}{2}\log(2\pi)}_{\text{Constant Term}}$ (21)
We optimized Equation 21 using L-BFGS to find:
  • Optimal task embeddings $\mathbf{e}_{\mathcal{S}}$ and $\mathbf{e}_{\mathcal{T}}$
  • Hierarchical distribution parameters $\bm{\mu}_{\text{e}}$ and $\sigma^{2}_{\text{e}}$
  • Other model hyperparameters (kernel parameters, noise variances, etc.)
The optimization balanced fitting the data well (minimizing the NLML term) while keeping task embeddings plausible under their common prior (minimizing the regularization term). The regularization strength adapted based on the learned variance $\sigma^{2}_{\text{e}}$.
This approach offered advantages for volatility construction. It provided a data-driven way to determine information transfer without fixed assumptions. When SABR structure matches market patterns well, embeddings could move closer, increasing information flow. When differences exist, embeddings could maintain separation while still benefiting from shared characteristics encoded by the kernel.
3.5 Prediction
Once we optimized all model parameters, we could make predictions for the target implied volatility surface at any new test point. The prediction followed standard GP principles, adapted for our multi-task framework.
Given optimized parameters $\hat{\bm{\theta}}$ and the combined training dataset $\mathcal{D}=\mathcal{D}_{\mathcal{S}}\cup\mathcal{D}_{\mathcal{T}}$, we computed the predictive distribution for the target function $f_{\mathcal{T}}$ at a new test point $\mathbf{x}_{\ast}=[K_{\ast},\tau_{\ast}]^{\intercal}$. This involved the joint Gaussian distribution of the target function value $f_{\mathcal{T}}(\mathbf{x}_{\ast})$ and all training observations $\mathbf{y}$:
$\begin{bmatrix}\mathbf{y}\\ f_{\mathcal{T}}(\mathbf{x}_{\ast})\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mathbf{0}_{M+N}\\ 0\end{bmatrix},\begin{bmatrix}\mathbf{K}\circ\mathbf{C}+\bm{\Sigma}&\mathbf{k}_{\ast}\\ \mathbf{k}_{\ast}^{\intercal}&C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\ast})\end{bmatrix}\right)$ (22)
The cross-covariance vector $\mathbf{k}_{\ast}$ related the test point to all training points across both tasks, incorporating learned task relationships via $C_{\mathcal{T,S}}$ and $C_{\mathcal{T,T}}$:
$\mathbf{k}_{\ast}=\Big{[}\underbrace{C_{\mathcal{T,S}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{S},1}),\dots,C_{\mathcal{T,S}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{S},M})}_{\text{Covariance with source (SABR) points}},\underbrace{C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{T},1}),\dots,C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{T},N})}_{\text{Covariance with target (market) points}}\Big{]}^{\intercal}$ (23)
The factor $C_{\mathcal{T,S}}$ controlled the influence of synthetic SABR data, while $C_{\mathcal{T,T}}$ determined the weight of the market data. These factors reflected the learned task relationship.
Using standard GP conditioning formulas, the predictive distribution for the target task at $\mathbf{x}_{\ast}$ was Gaussian:
$f_{\mathcal{T}}(\mathbf{x}_{\ast})|\mathcal{D}$ $\sim\mathcal{N}(\mu_{\ast},\sigma_{\ast}^{2})$ (24)
$\mu_{\ast}$ $=\mathbf{k}_{\ast}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{y}$ (25)
$\sigma_{\ast}^{2}$ $=C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\ast})-\mathbf{k}_{\ast}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{k}_{\ast}$ (26)
The mean $\mu_{\ast}$ was our estimation of implied volatility at $\mathbf{x}^{\ast}$, and the variance $\sigma_{\ast}^{2}$ quantified prediction uncertainty.
4 Experiments Design
To evaluate SABR-MTGP against benchmark methods, we designed numerical experiments. Since the goal was to assess implied volatility surface construction accuracy, we needed a reliable ground truth reflecting realistic market behavior. We chose the Heston stochastic volatility model [12] for this purpose, as it generates complex volatility dynamics similar to real markets.
Our experimental setup included generating realistic ground truth data, creating sparse “market" observations, calibrating the SABR model, training the different models, and evaluating their performance across market conditions. This section details the methodology; Section 5 presents the findings.
4.1 Data Generation Framework
Our experiments used two main datasets: (1) a sparse “market" dataset simulating real option price observations, and (2) a denser synthetic dataset derived from fitting a SABR model to these observations. The first was the target task, the second was the source task for multi-task learning.
4.1.1 Heston Ground Truth Data
We generated synthetic “market" data using the Heston model [12], which describes asset price $S_{t}$ and variance $v_{t}$ dynamics via coupled SDEs:
$dS_{t}$ $=(r-q)S_{t}dt+\sqrt{v_{t}}S_{t}dW_{t}^{(1)}$ (27)
$dv_{t}$ $=\kappa(\theta-v_{t})dt+\nu_{\text{vol}}\sqrt{v_{t}}dW_{t}^{(2)}$ (28)
Here, $r$ is the risk-free rate, $q$ the dividend yield, $\kappa$ the mean-reversion speed, $\theta$ the long-run variance, and $\nu_{\text{vol}}$ the volatility-of-volatility. Brownian motions $W_{t}^{(1)}$ and $W_{t}^{(2)}$ have correlation $\rho$.
We chose the Heston model because it generates realistic volatility smiles/skews and allows efficient option pricing via its characteristic function for accurate ground truth calculation. Moreover, its parameters have financial interpretations, enabling simulation of various market regimes. For our baseline scenario (the Base case), we used parameters: $v_{0}=0.09$, $\theta=0.09$, $\kappa=1.0$, $\nu_{\text{vol}}=0.8$, and $\rho=-0.8$. We set $S_{0}=100$, $r=0.03$, and $q=0.01$.
To simulate “market" option data, we generated a dataset $\mathcal{D}_{\mathcal{T}}$ containing $N=166$ European call option contracts. We chose the maturities to match typical market distribution. This led to expirations about monthly up to 9 months (0.08 to 0.75 years). Then, we used sparser intervals close to quarterly steps up to 1.75 years. And we included annual expirations at 2.0, 2.5, and 3.0 years. For each maturity, we chose strike prices ($K$) to cover a specific moneyness range ($K/S_{0}$). This range depended on the maturity: $(0.7,1.6)$ for short-term ($\tau\leq 0.5$ years), and $(0.8,1.4)$ for mid-term ($0.5<\tau\leq 1.5$ years) and long-term ($\tau>1.5$ years). Inside these ranges, we changed the strike intervals in a planned way. So, there were more strikes near the at-the-money (ATM) region (from $0.95S_{0}$ to $1.05S_{0}$) and fewer strikes further away. Specifically, the strike intervals around ATM, near ATM, and far from the money were (2.5, 5.0, 15.0) for short-term, (5.0, 10.0, 25.0) for mid-term, and (10.0, 25.0, 50.0) for long-term maturities. Figure 1 showed the data point distribution for our experiments.
Figure 1: Distribution of generated Heston “market" data points (target task $\mathcal{D}_{\mathcal{T}}$) in our experiments, by strike price and time to maturity. Dashed red lines indicate evaluation maturities.
For each contract, we computed the Heston price via Carr-Madan fast Fourier transform method [32]. Then, we obtained the Black-Scholes implied volatility $\sigma_{\text{mkt}}(K,\tau)$ by inverting the Heston price using a root-finding algorithm. This produced the “market" dataset $\mathcal{D}_{\mathcal{T}}=\{(\mathbf{X}_{\mathcal{T}},\mathbf{y}_{\mathcal{T}})\}$, the target task data.
4.1.2 SABR-Derived Synthetic Data
The SABR-MTGP used a denser synthetic dataset derived by fitting the SABR model to the Heston data. This involved several steps. First, for each unique maturity $\tau_{(p)}\in\mathbb{T}_{\text{mkt}}$, SABR parameters ($\alpha_{(p)},\rho_{(p)},\nu_{(p)}$) were calibrated by minimizing MSE between SABR volatilities $\sigma_{\text{SABR}}$ and market volatilities $\sigma_{\text{mkt}}$ from $\mathcal{D}_{\mathcal{T}}$ ($\beta=0.5$ fixed). We used Nelder-Mead optimization with constraints, yielding calibrated parameters $\{(\alpha_{(p)},\rho_{(p)},\nu_{(p)})\}$ for the set $\mathbb{T}_{\text{mkt}}$.
Second, we obtained SABR parameters for any time-to-maturity $\tilde{\tau}$ via piecewise linear interpolation (Equation 9) between parameters at the nearest surrounding calibrated maturities in $\mathbb{T}_{\text{mkt}}$. Constant extrapolation was used outside this range. Figure 2 shows the calibrated SABR parameters across maturities for the Base Heston scenario. The parameters showed term structures consistent with market behavior, supporting the interpolation.
Figure 2: Evolution of calibrated SABR parameters ($\alpha,\beta,\rho,\nu$) across maturity slices for the Base Heston scenario. Note $\beta$ was fixed at 0.5 during calibration.
Third, we defined a dense $35\times 35$ grid $(\tilde{K}_{j},\tilde{\tau}_{j})$, covering strike range ($K\in[70,160]$) and maturity range ( $\tau\in[0.08,3.0]$). At each grid point, we used the interpolated/extrapolated SABR parameters with the Hagan formula (Equation 7) to compute synthetic implied volatility $y_{\mathcal{S},j}=\sigma_{\text{SABR}}(\tilde{K}_{j},F_{\tilde{\tau}_{j}},\tilde{\tau}_{j};\alpha_{\tilde{\tau}_{j}},\beta,\rho_{{\tilde{\tau}}_{j}},\nu_{{\tilde{\tau}}_{j}}).$
Finally, we added small Gaussian noise $\epsilon_{\mathcal{S}}\sim\mathcal{N}(0,0.01^{2})$ to the synthetic data. This generated the synthetic source dataset $\mathcal{D}_{\mathcal{S}}=\{(\mathbf{X}_{\mathcal{S}},\mathbf{y}_{\mathcal{S}})\}$.
4.2 Comparative Models and Training
We compared three approaches for constructing the IVS. First, SABR Interpolation, a parametric baseline using calibrated and interpolated SABR parameters with the Hagan formula (Equation 7). Second, a Single-Task Gaussian Process (GP) trained only on the sparse Heston “market" data $\mathcal{D}_{\mathcal{T}}$. This GP used a Matérn 5/2 kernel on $(K,\tau)$ inputs and learned hyperparameters via L-BFGS optimization of marginal likelihood. Third, the proposed SABR-Informed Multi-Task GP (SABR-MTGP), our transfer learning framework. It was trained jointly on synthetic source data $\mathcal{D}_{\mathcal{S}}$ and market target data $\mathcal{D}_{\mathcal{T}}$, using the ICM structure with a shared Matérn 5/2 input kernel and a task kernel based on learned one-dimensional embeddings.
4.3 Evaluation Framework and Robustness Analysis
We evaluated model performance against the Heston ground truth on a predefined evaluation grid for fair comparison. This grid, separate from training data, included points $(K_{\ast},\tau_{\ast})$ spanning a relevant moneyness range (e.g., $K_{\ast}/S_{0}\in[0.8,1.4]$) at specific test maturities $\tau_{\ast}$ (e.g., 0.3, 0.9, 2.2 years) representing near-, mid-, and long-term. For each evaluation point, we computed the ground truth Heston IV using a high-precision FFT approach with the Heston characteristic function, followed by numerical inversion to obtain corresponding implied volatility.
Each model (SABR Interpolation, GP, SABR-MTGP) predicted $\hat{\sigma}(K_{\ast},\tau_{\ast})$ on this grid. We measured performance using Root Mean Squared Error (RMSE) and Mean Absolute Error (MeanAE) against the ground truth $\sigma_{\text{Heston}}(K_{\ast},\tau_{\ast})$, calculated over valid predictions.
To assess robustness, we repeated experiments using ten Heston parameter configurations (Table 1), including the Base scenario. These settings simulated various market dynamics by altering parameters like mean reversion $\kappa$, volatility levels $\theta,\nu_{\text{vol}}$, correlation $\rho$, and initial term structure $v_{0}$ vs $\theta$. This allowed analysis across different market conditions. Robustness evaluation focused on metrics at distinct test maturities. We also analyzed the MTGP’s learned task correlation and variance decomposition for each scenario to understand its adaptive information-sharing.
Table 1: Heston Parameter Configurations for Robustness Analysis. Common parameters are $S_{0}=100,r=0.03,q=0.01$.
Setting Name $\kappa$ $\theta$ $\nu_{\text{vol}}$ $\rho$ $v_{0}$
Base 1.0 0.09 0.8 -0.8 0.09
Moderate Mean-Rev 2.0 0.09 0.8 -0.8 0.09
Low Mean-Rev 0.5 0.09 0.8 -0.8 0.09
High Vol-Regime 1.0 0.16 0.9 -0.8 0.16
Low Vol-Regime 1.0 0.04 0.4 -0.8 0.04
Moderate Correlation 1.0 0.09 0.8 -0.5 0.09
Strong Correlation 1.0 0.09 0.8 -0.9 0.09
Term Structure Up 1.0 0.16 0.8 -0.8 0.09
Term Structure Down 1.0 0.04 0.8 -0.8 0.16
Mixed Regime 1.5 0.12 0.6 -0.6 0.12
5 Results and Discussion
This section presents our findings on SABR-MTGP performance compared with SABR interpolation and the standard GP. We first analyze performance under the Base Heston scenario, then examine robustness across various market conditions.
5.1 Performance in the Base Scenario
Figures 3(a), 3(b), and 3(c) show fitted implied volatility curves from the three models compared against Heston ground truth, for near-term ($\tau=0.30$), mid-term ($\tau=0.90$), and long-term ($\tau=2.20$) maturities in the Base scenario. Each figure includes fitted curves and residuals.
(a) Near-term ($\tau=0.30$).
(b) Mid-term ($\tau=0.90$).
(c) Long-term ($\tau=2.20$).
Figure 3: Implied volatility curve comparison and residuals under the Base Heston scenario for (a) Near-term, (b) Mid-term, and (c) Long-term maturities.
The results showed different performance patterns across the term structure. For near-term maturity (Figure 3(a)), where market data is relatively dense, SABR-MTGP achieved the most accurate fit across strikes (RMSE=0.0014). The standard GP (RMSE=0.0024) showed systematic deviations, suggesting difficulty capturing the precise smile pattern. Direct SABR interpolation (RMSE=0.0057) struggled with curvature, showing systematic errors consistent with limitations of its parametric form in matching these Heston dynamics.
Moving to mid-term maturity (Figure 3(b)), data sparsity increased. SABR-MTGP continued to provide the closest fit (RMSE=0.0005). SABR interpolation also performed very well (RMSE=0.0009), capturing the overall shape effectively. However, the standard GP (RMSE=0.0043) showed larger and more scattered errors, particularly for strikes away from the money.
For long-term maturity (Figure 3(c)), SABR interpolation performed best according to RMSE (0.0005), benefiting from its structure to provide a reasonable approximation. SABR-MTGP also achieved very good accuracy (RMSE=0.0007). The standard GP (RMSE=0.0025) exhibited notably higher errors compared to the other two methods, with residuals spanning approximately from $-0.006$ to $+0.002$ volatility points, indicating difficulties in generalizing from the sparse long-term data.
These findings highlight the advantages of our hybrid approach. In data-rich regions (near-term), SABR-MTGP combined market data with structural guidance to outperform alternatives. As data gets sparser (mid- to long-term), it used SABR structure effectively while retaining flexibility to adapt to Heston dynamics, balancing SABR’s rigidity and standard GP’s data dependence.
5.2 Robustness Across Market Conditions
The performance heatmaps in Figures 4, 5, and 6 summarized model accuracy (RMSE and MeanAE) across the ten Heston parameter settings, evaluated at near-, mid-, and long-term horizons. Lighter shades indicated lower errors.
Figure 4: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Near-term maturity ($\tau=0.3$).
Figure 5: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Mid-term maturity ($\tau=0.9$).
Figure 6: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Long-term maturity ($\tau=2.2$).
These heatmaps show consistent performance patterns. SABR-MTGP consistently demonstrated strong performance, frequently ranking as the best or a very close second-best method. The standard GP showed reasonable performance in the near-term, but its accuracy dropped significantly at long-term maturities, where it yielded the highest errors. Conversely, SABR interpolation struggled in the near-term, often producing the highest errors, but benefited from its fixed structure in long-term cases. It frequently achieved the lowest errors, especially with very sparse data. SABR-MTGP’s ability to adaptively blend structural guidance with data-driven flexibility allowed it to maintain high accuracy overall, overcoming SABR’s near-term issues and the GP’s long-term instability.
5.3 Learned Task Relationship
The MTGP framework learned the relationship between the SABR source task and the Heston market target task. Figure 7 shows this adaptive relationship for three Heston settings, illustrating variations in learned correlation, covariance structure, and variance decomposition (task-specific proportion $\kappa^{2}_{\mathcal{Z}}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$ vs. shared variance proportion $\sigma^{2}_{h}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$).
(a) Low Correlation Example: Moderate Mean-Rev Setting (Learned Correlation = 0.22).
(b) Moderate Correlation Example: Base Setting (Learned Correlation = 0.41).
(c) High Correlation Example: Term Structure Up Setting (Learned Correlation = 0.90).
Figure 7: Examples of learned task relationship analysis for the SABR-MTGP model under different Heston parameter settings: (a) Moderate Mean-Rev, (b) Base, and (c) Term Structure Up. Panels show variations in learned correlation (left), covariance structure (center), and variance decomposition (right).
The learned correlation varied with Heston parameters. Faster mean reversion (Moderate Mean-Rev) yielded low correlation (0.22), perhaps because SABR captured these dynamics less well. An upward term structure (Term Structure Up) resulted in high correlation (0.90), suggesting SABR provided a better structural approximation in this case. The Base setting showed moderate correlation (0.41). This showed the model’s ability to assess similarity between SABR information and target data based on market conditions. The task covariance matrices reflected these correlations.
A key pattern was that the shared covariance component consistently accounted for most variance in both tasks across scenarios. Even if SABR is not a perfect match (low correlation), the MTGP leveraged shared properties like smoothness, preventing overfitting to sparse target data and performing better than the standard GP at longer maturities. When SABR aligns well (high correlation), the model strongly used this information, achieving accuracy competitive with or better than SABR interpolation, while maintaining flexibility.
5.4 Parameter Sensitivity Analysis
We further examined robustness by analyzing performance sensitivity to Heston parameter changes. Using the coefficient of variation (CV) of RMSE, Figure 8 showed how much the precision of each model fluctuated within the parameter categories relative to its average performance. Lower CV (lighter colors) indicated more stable performance.
(a) Near-term ($\tau=0.3$).
(b) Mid-term ($\tau=0.9$).
(c) Long-term ($\tau=2.2$).
Figure 8: Parameter sensitivity heatmaps showing the Coefficient of Variation (CV) of RMSE for GP, SABR, and SABR-MTGP across different Heston parameter categories (y-axis) for near-, mid-, and long-term evaluation maturities. Lower CV (lighter color) indicated greater performance stability within that parameter category.
The sensitivity heatmaps highlighted model stability differences. The standard GP often showed high CV, indicating its performance was sensitive to underlying volatility dynamics, especially concerning ‘Correlation’ and ‘Vol-of-Vol’ changes at longer maturities. SABR interpolation typically had the lowest CV due to its fixed parametric form, but this stability could come at the cost of lower accuracy, particularly near-term. SABR-MTGP generally displayed intermediate stability. It demonstrated improved stability over the standard GP in the mid- and long-term. Although it was less stable than SABR interpolation. The sensitivity results thus complemented accuracy findings: SABR-MTGP found a middle ground, using the prior for improved stability over GP, while retaining adaptive flexibility unlike SABR interpolation.
5.5 Summary
The experiments using Heston data suggested SABR-MTGP effectively constructed the IVS. By combining SABR’s structural information with market observations in a multi-task framework, it achieved better accuracy and stability than using either a standard GP or SABR. It effectively integrated structural information with observed data, adaptively leveraging SABR insights by assessing their relevance to the target dataset.
6 Application to SPX Market Data
To further evaluate the practical applicability of SABR-MTGP, we tested its performance on real-world market data for SPX options (provided by the OptionMetrics database), observed on August 1, 2023. This case study complements the controlled experiments using the Heston model by assessing the method’s behavior under actual market conditions.
6.1 Data
The dataset included European call options on the SPX index. We preprocessed the data by filtering for valid quotes, calculating time-to-maturity, and finding implied volatilities from mid-prices. We used the Zero Coupon Bond yield curve as the interest rate curve, and estimated the dividend yield using put-call parity. This gave us several hundred valid market implied volatility observations across different strikes and maturities up to about one year. Key statistics summarizing this dataset are in Table 2. This processed market data was the target task ($\mathcal{D}_{\mathcal{T}}$) for both the SABR-MTGP model.
Table 2: Descriptive Statistics for the Preprocessed SPX Call Option Dataset (August 1, 2023)
Statistic Value
Total Options 524
Unique Maturities 12
Maturity Range (Years) [0.0465, 0.9665]
Strike Price Range [3300.00, 6600.00]
Implied Volatility Range [0.0986, 0.3322]
Moneyness Distribution (Based on $K/S_{0}$, $S_{0}=4576.73$)
In-the-Money ($K/S_{0}<0.95$) 68 (13.0%)
At-the-Money ($0.95\leq K/S_{0}\leq 1.05$) 289 (55.2%)
Out-of-the-Money ($K/S_{0}>1.05$) 167 (31.9%)
  • Note: Statistics derived from the preprocessed SPX dataset used in this study. The moneyness definition uses $S_{0}=4576.73$.
6.2 Results
We do not have a ground truth IVS for real market data. So, the evaluation used a qualitative assessment of the model fits and the resulting surfaces. Figure 9 shows typical implied volatility slices for near-term ($\tau=0.2190$ years) and mid-term ($\tau=0.8898$ years) maturities.
(a) Near-term maturity ($\tau=0.2190$ years).
(b) Mid-term maturity ($\tau=0.8898$ years).
Figure 9: Implied volatility slices for SPX options on August 1, 2023. Comparison of SABR-MTGP, GP, and SABR interpolation against market data points. Residuals (Model - Market) are shown below each slice.
In both slices (Figure 9), all three methods generally captured the observed smile/skew pattern of the market. But the residual plots showed important differences in fit quality. For the near-term slice (Figure 9(a)), SABR interpolation had systematic errors. It underestimated volatility for low strikes (ITM) and overestimated near the money. This was shown by the consistent pattern in its residuals. Both the standard GP and SABR-MTGP fitted the observed market points better, especially around the at-the-money region. But there was a notable difference in the deep in-the-money region (strikes 3250-4000). The standard GP curve, although sloping downward, showed a shape considered less regular and potentially less financially realistic compared to volatility curve given by the SABR. The SABR-MTGP curve maintained a smoother, decreasing volatility in this data-sparse region. This produced a more financially realistic interpolation, likely because of the structural guidance from the SABR source task. For the mid-term slice (Figure 9(b)), where the smile was flatter and the market data was denser, all models achieved a very good fit. Notably, the SABR model fit improved significantly compared to the near-term slice. The residuals for both GP and SABR-MTGP were particularly small (generally within +/- 0.001), demonstrating excellent agreement with market observations.
(a) GP
(b) SABR Interpolation
(c) SABR-MTGP
Figure 10: SPX Implied Volatility Surfaces constructed using (a) GP, (b) SABR Interpolation, and (c) SABR-MTGP, based on market data from August 1, 2023. Black dots represent observed market implied volatilities.
The three plots (Figure 10) showed the different shapes of the implied volatility surfaces produced by the three models. The standard GP surface (Figure 10(a)) matched the observed market data points very closely. However, in data-sparse regions, particularly for deep out-the-money options ($K$ is large), it showed some unusual patterns that were inconsistent with typical market behavior, such as a downward slope in volatility where an upward sloping skew is generally expected. The SABR surface (Figure 10(b)) was naturally smooth because it was based on a specific mathematical formula. But, because of this fixed structure, it might not have captured all the specific details or patterns seen in the market data. The SABR-MTGP surface (Figure 10(c)) seemed to offer a good balance between the other two methods. It fitted the market data well, similar to the standard GP. And, it maintained a smoother and more economically consistent shape compared to the standard GP, especially when extending to areas with less data, such as longer times-to-maturity or strike prices far from the current price. This visual comparison suggested that the SABR-MTGP model successfully used the general structure of the SABR model to guide the construction of the surface, particularly where market data was sparse. This helped to create a more stable and realistic surface overall.
In summary, the application to real SPX market data showed the model’s ability to integrate structural financial knowledge with sparse empirical observations. This led to well-behaved and plausible implied volatility surfaces suitable for practical applications.
7 Conclusion
In this paper, we proposed SABR-MTGP, an approach bridging structural and data-driven methods for IVS construction. The key idea is framing IVS construction as a multi-task learning problem where we used structural information from the SABR model to improve predictions from sparse market data.
Our approach offered several contributions. First, we show how to combine financial models with machine learning techniques. We used SABR as a complementary information source within our multi-task framework. Second, we developed a hierarchical Bayesian regularization for task embeddings that adaptively determined the level of information transfer. Third, our evaluation using Heston ground truth data showed this approach balanced structure and flexibility. It provided reliable predictions across market conditions and outperformed standard Gaussian process particularly in data-sparse regions. Improvements were notable for longer maturities and extreme strikes, where market data was limited. Finally, an application to real SPX market data further illustrated the model’s ability to produce smooth, reliable IVS that respect market observations while benefiting from structural guidance.
Our work has practical implications for quantitative finance. SABR-MTGP provides a more reliable tool for IVS construction, crucial for derivative pricing, hedging, and risk management. By incorporating structural knowledge and empirical observations, it reduced model misspecification risk while maintaining adaptability to market realities. Future research could extend the multi-task framework to other structural models or multiple models simultaneously.

### v2

@@
 Abstract
-Constructing the Implied Volatility Surface (IVS) is a challenging task in quantitative finance due to the complexity of real markets and the sparsity of market data.
-Structural models like Stochastic Alpha Beta Rho (SABR) model offer interpretability and theoretical consistency but lack flexibility, while purely data-driven methods such as Gaussian Process regression can struggle with sparse data.
-We introduce SABR-Informed Multi-Task Gaussian Process (SABR-MTGP), treating IVS construction as a multi-task learning problem.
-Our method uses a dense synthetic dataset from a calibrated SABR model as a source task to inform the construction based on sparse market data (the target task).
-The MTGP framework captures task correlation and transfers structural information adaptively, improving predictions particularly in data-scarce regions.
-Experiments using Heston-generated ground truth data under various market conditions show that SABR-MTGP outperforms both standard Gaussian process regression and SABR across different maturities.
-Furthermore, an application to real SPX market data demonstrates the method’s practical applicability and its ability to produce stable and realistic surfaces.
-This confirms our method balances structural guidance from SABR with the flexibility needed for market data.
+This study introduces a SABR-informed multitask Gaussian process for constructing implied volatility surfaces from sparse option quotes.
+We treat a dense synthetic dataset generated by a calibrated SABR model as the source task and market option quotes as the target task.
+Within the multitask Gaussian process framework, we learn cross-task dependence via task embeddings with hierarchical regularization, enabling adaptive transfer of structural information.
+On Heston ground truth across ten market regimes and in a case study with SPX options, the model achieves lower error than the single-task Gaussian process and SABR at near-term maturities and remains competitive at long-term maturities, while satisfying standard no-arbitrage conditions.
+The framework combines the theory-driven structure with nonparametric Bayesian regression and yields reliable implied volatility surfaces for risk management.
 keywords:
-Gaussian Process , Implied Volatility , Multi-Task Learning , SABR , Machine Learning.
+Gaussian Process , Implied Volatility , Multitask Learning , SABR , Machine Learning
+journal: Journal of Computational Science
 […]
@@
 1 Introduction
-Constructing the Implied Volatility Surface (IVS) is an important and difficult task in quantitative finance.
-A common difficulty arises from sparse market option data, especially for long-dated expiration or strikes far from the money.
-This sparsity causes difficulties for conventional construction approaches.
+Constructing the implied volatility surface (IVS) from sparse option quotes is a standard task in practice.
+Trading desks use the surface to support volatility quoting and the day-to-day risk management of vanilla option books across strikes and maturities.
+Structuring teams and model developers use it as an input to the calibration of downstream pricing models.
+The task is particularly challenging when market liquidity is limited, for example around the market open or for less liquid option products, and in illiquid regions such as far-from-the-money strikes and long maturities where observations are sparse and unevenly distributed.
+These operational requirements motivate surface construction methods that remain stable under data sparsity while producing smooth and no-arbitrage extrapolations in the wings.
 Structural models, such as the SABR model [23], are popular because they capture volatility smiles and skews with few parameters.
-The SABR model assume that the underlying asset and its volatility follow specific stochastic differential equations.
-While these models provide reasonable interpolations, their predetermined functional forms may suffer from a lack of flexibility and not fully reflect market complexities.
-In contrast, data-driven approaches offer more flexibility, adapting to market patterns without imposing rigid structures.
-Gaussian Processes (GP) [2], provide a Bayesian framework for regression.
-However, these methods typically require substantial data to achieve reliable results.
-With sparse observations, GP may fail to capture patterns of implied volatility.
-In order to address the challenges in data-driven approaches, recent works discuss how to incorporate machine learning methods with prior financial knowledge.
-Cousin et al.
-[3], Chataigner et al.
-[4], Ackerer et al.
-[5], Zheng et al.
-[6], Gonon et al.
-[7], Hoshisashi et al.
-[8] present methods that imposing constraints derived from financial theory within a neural network or a Gaussian Process.
+These models assume that the underlying asset and its volatility follow specific stochastic differential equations.
+While these models provide reasonable interpolations, their predetermined functional forms may lack the flexibility to fully reflect market complexities.
+In contrast, data-driven approaches such as the Gaussian process (GP) [38], a nonparametric Bayesian method that can be used in IVS construction [15, 36], offer more flexibility.
+However, this adaptability requires sufficient data.
+With sparse observations, they can be prone to overfitting and may violate no-arbitrage conditions.
+In order to address the challenges in data-driven approaches, recent works discuss how to incorporate prior financial knowledge into machine learning methods.
+The literature [12, 7, 1, 41, 20, 27] presents methods that impose constraints derived from financial theory within a neural network or a Gaussian process.
 In contrast, Chen et al.
-[9] use typical transfer learning methods for option pricing.
-A neural network is trained on a structural model’s synthetic data (generated by the Black-Scholes model) is primarily, and it provide a good initialization for subsequent fine-tuning on empirical data.
-Different to their methods, we present the SABR-Informed Multi-Task Gaussian Process (SABR-MTGP) which takes advantage of both structural and data-driven models.
+[8] use transfer learning methods for option pricing.
+A neural network is trained on synthetic data generated from the Black-Scholes model, and it provides a good initialization for subsequent fine-tuning on empirical data.
+Inspired by the transfer learning approach of Chen et al.
+[8], we propose a SABR-informed multitask Gaussian process (SABR-MTGP).
 The key idea is to treat the SABR-generated IVS as an information source (a ‘source’ task) and market observations as the primary target task.
-Then, a multi-task learning framework [10, 11] is used where the relationship between the theoretical structure (SABR) and empirical observations (market data) is learned adaptively during a joint optimization process.
+Then, a multitask learning framework [5, 2] is used where the relationship between the theoretical structure (SABR) and empirical observations is learned adaptively during a joint optimization process.
+In our experiments, the proposed SABR-MTGP stabilizes data-sparse wings (far-from-the-money strikes) relative to a single-task GP, while correcting systematic misspecification that can arise from purely SABR models.
 The implementation consists of two main stages.
 First, we generate a synthetic dataset using the calibrated SABR model.
-This dataset embodies smile and term-structure characteristics typical of SABR dynamics.
-Second, we train a multi-task Gaussian Process (MTGP) model simultaneously on these dense synthetic data and the sparse market observations.
-The MTGP framework learns correlation patterns between tasks through shared and task-specific covariance components.
-We use task embeddings with hierarchical regularization to appropriately balance how much structural guidance from SABR is used when constructing the representation from real market data.
-This avoids both overly rigid adherence to SABR patterns and complete disregard for its structural insights.
-We evaluated our method through numerical experiments using the Heston model [12] to generate sparse ‘market’ datasets.
-Comparing SABR-MTGP against standard GP and SABR with interpolation, we find that our approach generally yields more accurate volatility surfaces across diverse market conditions and maturities, with particular advantages in sparse data regions.
-We further demonstrate the practical applicability of our method through a case study on real SPX market data.
-These results show the effectiveness of our method in integrating financial domain knowledge within a flexible machine learning framework.
+This dataset embodies smile and term-structure patterns characteristic of SABR dynamics.
+Second, we train a multitask Gaussian process (MTGP) model simultaneously on these dense synthetic data and the sparse market observations.
+The MTGP framework learns the correlation between tasks through the shared covariance component.
+We use task embeddings with hierarchical regularization to appropriately balance how much structural guidance from SABR is used when constructing the IVS from real market data.
+SABR-MTGP improves IVS reconstruction from sparse quotes, but it is not a replacement for SABR-based dynamic hedging, or OTC derivatives valuation.
+Our scope is IVS construction from sparse observations.
+Extending to hedging and OTC derivatives valuation requires market-consistent dynamics (e.g., local or stochastic local volatility) and Greek validation.
+These are left to future work.
 Related Literature
-Gaussian processes have found increasing application across finance and economics due to their flexibility and ability to quantify uncertainty.
+Implied volatility surface construction can be organized into three complementary approaches.
+First, structural models specify dynamics that reproduce smiles and term structures with interpretable parameters, including SABR [23], Heston [25], Variance Gamma [32] and the double-exponential model [28].
+More recently, rough volatility models [3, 18] emphasize fractional regularity in volatility paths.
+Second, parametric surface parameterizations impose shape and no-arbitrage structure, e.g., SVI [19], SSVI [17], and spline-based families [11].
+Third, machine learning approaches, including the Gaussian process [15, 36, 7] and neural networks [1, 41, 20, 27, 34, 10] provide data-driven alternatives.
+Relative to transfer learning approaches for options (Chen et al.
+[8]), our contribution is to implement the transfer learning in the Gaussian process via a task kernel with learnable embeddings and a hierarchical prior that adapts cross-task correlation.
+Besides implied volatility surface construction, Gaussian processes have found increasing application across finance and economics due to their flexibility and ability to quantify uncertainty.
 A significant body of work focuses on derivative pricing, valuation, and hedging.
-GP models have been used for constructing implied volatility curves [13, 14], constructing financial term structures [15], accelerating pricing and hedging calculations for various options [13, 16, 17, 18, 19], modeling derivative portfolios for CVA computation [20], pricing complex insurance products like variable annuities [21], and approximating Greeks for hedging [22].
+GP models have been used for constructing financial term structures [13], accelerating pricing and hedging calculations for various options [15, 31, 21, 29, 26], modeling derivative portfolios for CVA computation [14], pricing complex insurance products like variable annuities [22], and approximating Greeks for hedging [30].
 Beyond direct pricing, GP models serve as flexible tools for calibrating implied and local volatility surfaces [36, 7] or measuring portfolio tail risk [35].
 Furthermore, GP models are employed in broader financial and economic time series analysis, including imputing missing financial data [16], forecasting real estate prices [39], modeling inflation dynamics [9], understanding determinants of carbon market prices [37], and developing nonparametric vector autoregressions for macroeconomic analysis [24].
-While these studies highlight the versatility of GP models, our work specifically contributes by employing a multi-task learning framework informed by a structural model (SABR) to address the challenge of sparse data in IVS construction, learning the relationship between the structural prior and market data adaptively.
-The remainder of the paper is organized as follows: Section 2 introduces some basic concept in implied volatility surface and Gaussian Process.
+The remainder of the paper is organized as follows: Section 2 introduces basic concepts of the implied volatility surface.
 Section 3 provides the details of the proposed methods.
 In Section 4, we describe the setup of the numerical experiments and the results are presented in Section 5.
-In Section 6, we demonstrate the performance of the proposed method in real market data.
-2 Background
-This section provides a necessary background on the implied volatility surface and the Gaussian process.
-2.1 Implied Volatility Surface
-The Black-Scholes model [29] is a basic framework for European option pricing under specific assumptions: key assumptons include that the underlying asset price follows the Geometric Brownian Motion, and asset volatility are constants.
-Observed market option prices frequently deviate from the prices given by the Black-Scholes model under the assumption of constant volatility.
-This difference is captured through the Implied Volatility (IV), a way to match the theoretical framework with market observations.
-For an observed market price of call options $C_{mkt}(K,\tau)$ with time-to-maturity $\tau$ and strike price $K$, the corresponding IV $\sigma_{mkt}(K,\tau)$ is the unique positive value that makes the following equation holds:
+In Section 6, we demonstrate the application of the proposed methods using real market data.
+2 Background: Implied Volatility Surface
+The Black-Scholes model [4] is a basic framework for European option pricing under specific assumptions: the underlying asset price follows a geometric Brownian motion, and the asset’s volatility is constant.
+In practice, this constant volatility assumption is inconsistent with observed option prices, which typically exhibit patterns such as volatility smiles and skews.
+To reconcile the Black-Scholes formula with market prices, the concept of Implied Volatility (IV) is introduced.
+For an observed market price of call options $C_{mkt}(K,\tau)$ with time-to-maturity $\tau$ and strike price $K$, the corresponding IV $\sigma_{mkt}(K,\tau)$ is the unique positive value that makes the following equation hold:
 $C_{mkt}(K,\tau)=C_{BS}(K,\tau,\sigma=\sigma_{mkt}(K,\tau)\,;\theta_{BS}),$ (1)
-where $C_{BS}$ is the Black-Scholes formula, and $\theta_{BS}$ is the collection of rest parameters in the Black-Scholes formula.
-Finding $\sigma_{mkt}(K,\tau)$ needs solving the Black-Scholes formula backwards, usually done with iterative root-finding algorithms (like Newton-Raphson).
+where $C_{BS}$ is the Black-Scholes formula, and $\theta_{BS}$ is the collection of remaining parameters in the Black-Scholes formula.
+Computing $\sigma_{mkt}(K,\tau)$ requires inverting the Black-Scholes formula using iterative root-finding algorithms (such as Newton-Raphson).
 When $\sigma_{mkt}(K,\tau)$ is shown as a function of both variables, the resulting three-dimensional surface is the Implied Volatility Surface (IVS).
@@
 Accurate IVS construction is important for derivative pricing, hedging strategies, and risk management across option portfolios.
-2.2 Gaussian Process
-A Gaussian Process (GP) is a Bayesian non-parametric approach that can be used in regression problems [2].
-A GP defines a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$ such that any finite set of function values has a joint Gaussian distribution.
-A GP is fully specified by its mean function $m:\mathbb{R}^{d}\to\mathbb{R}$ and covariance function (or kernel) $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$:
-$f\sim\mathcal{GP}(m,k).$
-The mean function is often set to zero for simplicity.
-The kernel encodes assumptions about function properties like smoothness or characteristic length-scales, defining the covariance between function values at any two input points.
-A commonly used kernel is the Matérn 5/2 kernel function which is a second-order differentiable kernel function:
-[…]
-$\text{for}\quad\mathbf{x,x^{\prime}}\in\mathbb{R}^{d}$, where $\sigma_{f}$ is the scaling hyperparameter and $\gamma$ is the length-scales hyperparameter.
-Given training data $\mathcal{D}=\{(\mathbf{X,y})\}$ where the components in $\mathbf{y}=[y_{1},\dots,y_{N}]^{\intercal}$ are noisy observations at $N$ locations $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{N}]^{\intercal}$.
-$y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ where $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{n}^{2})$ is noise.
-As a result, a predictive distribution for the function value $f_{\ast}$ at a new test point $\mathbf{x}_{\ast}$ is given by the posterior $p\big{(}f_{\ast}(\mathbf{x}_{\ast})|\mathcal{D}\big{)}\sim\mathcal{N}(\mu_{\ast},\sigma_{\ast})$, where
-$\mu_{\ast}$ $=k(\mathbf{X,x_{\ast}})^{\intercal}\big{(}k(\mathbf{X,X})+\sigma_{n}^{2}\mathbf{I}\big{)}^{-1}\mathbf{y},$ (2)
-$\sigma_{\ast}^{2}$ $=k(\mathbf{x}_{\ast},\mathbf{x}_{\ast})-k(\mathbf{X,x_{\ast}})^{\intercal}\big{(}k(\mathbf{X,X})+\sigma_{n}^{2}\mathbf{I}\big{)}^{-1}k(\mathbf{X,x_{\ast}}).$ (3)
-Here, $k(\mathbf{X,x_{\ast}})\coloneqq\left[k(\mathbf{x}_{1},\mathbf{x_{\ast}}),\dots,k(\mathbf{x}_{N},\mathbf{x_{\ast}})\right]^{\intercal}$, $\mathbf{I}$ is an identity matrix and
-$k(\mathbf{X,X})\coloneqq\begin{bmatrix}k(\mathbf{x}_{1},\mathbf{x}_{1})&\dots&k(\mathbf{x}_{1},\mathbf{x}_{N})\\ \vdots&\ddots&\vdots\\ k(\mathbf{x}_{N},\mathbf{x}_{1})&\dots&k(\mathbf{x}_{N},\mathbf{x}_{N})\end{bmatrix}.$
-The kernel hyperparameters and noise variance $\sigma_{n}^{2}$ are usually estimated by maximizing the marginal likelihood of the data.
-This provides an automatic way to balance data fit and model complexity [30].
-GP provide a Bayesian framework that can model complex, non-linear relationships without imposing rigid functional forms.
-However, they can face challenges when training data is sparse, as often happens with implied volatility observations.
-3 SABR-Informed Multi-Task Gaussian Process
+3 SABR-Informed Multitask Gaussian Process
 In this section, we describe our approach for constructing the implied volatility surface by combining SABR structure with GP flexibility.
 3.1 Problem Formulation
-The primary objective of this work is to construct an accurate Implied Volatility Surface (IVS).
-The IVS, denoted as $\sigma(K,\tau)$, is a function of the option strike price $K$ and its time-to-maturity $\tau$.
+The objective of this work is to construct an accurate Implied Volatility Surface (IVS).
+The IVS is a function of the option strike price $K$ and its time-to-maturity $\tau$.
 We can represent the input features as a two-dimensional vector $\mathbf{x}=[K,\tau]^{\intercal}$.
-The task is thus to learn a regression model $f:\mathbb{R}^{2}\to\mathbb{R}^{+}$ that maps these input features to the corresponding implied volatility, i.e., $y=f(\mathbf{x})=\sigma(K,\tau)$.
+The task is to learn a regression model $f:\mathbb{R}^{2}\to\mathbb{R}^{+}$ that maps these input features to the corresponding implied volatility.
 Suppose that we have $N$ observed locations $\mathbf{X}_{\mathcal{T}}=[\mathbf{x}_{\mathcal{T},1},\dots,\mathbf{x}_{\mathcal{T},N}]^{\intercal}$, $\mathbf{x}_{\mathcal{T},i}=[K_{i},\tau_{i}]^{\intercal}$, and $\mathbf{y}_{\mathcal{T}}=[y_{\mathcal{T},1},\dots,y_{\mathcal{T},N}]^{\intercal}$ contains the corresponding observed market implied volatilities, $y_{\mathcal{T},i}=\sigma_{\rm mkt}(K_{i},\tau_{i})$.
@@
 This synthetic dataset acts as a source of information to guide the learning process and prediction, especially in regions where market data is scarce.
-The subsequent sections will show how this synthetic data is generated and incorporated within a multi-task learning framework.
+The subsequent sections will show how this synthetic data is generated and incorporated within a multitask learning framework.
 3.2 Synthetic Data Generation via SABR
@@
 Hagan’s asymptotic expansion provides a closed-form approximation for implied volatility under SABR, which is crucial for efficiently generating the dense synthetic dataset $\mathcal{D}_{\mathcal{S}}$.
-The SABR model describes the dynamics of the forward price $F_{t}\coloneqq S_{t}e^{(r-q)(T-t)}$ (where $S_{0}$ is the current price of underlying asset, $r$ is the risk-free interest rate, $q$ is the dividend rate, $T$ is the maturity) and its instantaneous volatility $\alpha_{t}$ via stochastic differential equations:
+The SABR model describes the dynamics of the forward price $F_{t}\coloneqq S_{t}e^{(r-q)(T-t)}$ (where $S_{0}$ is the current price of the underlying asset, $r$ is the risk-free interest rate, $q$ is the dividend rate, $T$ is the maturity) and its instantaneous volatility $\alpha_{t}$ via stochastic differential equations:
 $dF_{t}$ $=\alpha_{t}F_{t}^{\beta}dW_{t}^{(1)},$ (2)
@@
 Practical applications often use Hagan’s asymptotic expansion [23], which provides closed-form approximations for implied volatility $\sigma_{\rm SABR}(K,\tau,F;\alpha,\beta,\rho,\nu)$.
-$\sigma_{\rm SABR}(K,\tau,F)=\left\{\begin{aligned} &I^{0}(K,\tau,F)\left[1+I^{1}(K,F)\tau\right]\quad&\text{if}\,K\neq F,\\ &\frac{\alpha}{F^{1-\beta}}\left[1+\left(\frac{(1-\beta)^{2}\alpha^{2}}{24F^{2-2\beta}}+\frac{\rho\beta\alpha\nu}{4F^{1-\beta}}+\frac{2-3\rho^{2}}{24}\nu^{2}\right)\tau\right]\quad&\text{if}\,K=F.\end{aligned}\right.$ (7)
+$\sigma_{\rm SABR}(K,\tau,F)=\left\{\begin{aligned} &I^{0}(K,\tau,F)\left[1+I^{1}(K,F)\tau\right]\quad&\text{if}\,K\neq F,\\ &\frac{\alpha}{F^{1-\beta}}\left[1+\left(\frac{(1-\beta)^{2}\alpha^{2}}{24F^{2-2\beta}}+\frac{\rho\beta\alpha\nu}{4F^{1-\beta}}+\frac{2-3\rho^{2}}{24}\nu^{2}\right)\tau\right]\quad&\text{if}\,K=F.\end{aligned}\right$.
+(5)
 where:
@@
 Let $\mathbb{T}_{\text{mkt}}=\{\tau_{(1)},\dots,\tau_{(P)}\}$ be the set of unique maturities in the market data $\mathcal{D}_{\mathcal{T}}$, ordered $\tau_{(1)}<\tau_{(2)}<\dots<\tau_{(P)}$.
-For each maturity slice $\tau_{(p)}$, we calibrated SABR parameters $(\alpha_{(p)},\rho_{(p)},\nu_{(p)})$ to match the observed market smile by minimizing the squared errors by Nelder-Mead algorithm [31]:
+For each maturity slice $\tau_{(p)}$, we calibrated SABR parameters $(\alpha_{(p)},\rho_{(p)},\nu_{(p)})$ to match the observed market smile by minimizing the squared errors using the Nelder-Mead algorithm [33]:
 $\min_{\alpha_{(p)},\rho_{(p)},\nu_{(p)}}\sum_{i:\,\tau_{i}=\tau_{(p)}}\left(\sigma_{\rm SABR}(K_{i},F_{\tau_{(p)}},\tau_{(p)};\,\alpha_{(p)},\beta,\rho_{(p)},\nu_{(p)})-y_{\mathcal{T},i}\right)^{2}$ (6)
 Here, $F_{\tau_{(p)}}=S_{0}e^{(r-q)\tau_{(p)}}$ is the forward price at maturity $\tau_{(p)}$.
-We fixed $\beta=0.5$ throughout our experiments, a common choice for equity options.
 Calibrating all four SABR parameters ($\alpha,\beta,\rho,\nu$) simultaneously for each maturity slice can be challenging and numerically unstable, especially with sparse or noisy data.
 Calibrating $\beta$ often requires richer data across strikes than might be available, particularly for longer maturities.
-Fixing $\beta$ reduces the dimensionality of the optimization problem for each slice from four parameters to three ($\alpha,\rho,\nu$), significantly improving the stability and speed of the calibration process.
-The parameters $\alpha_{(p)}$, $\rho_{(p)}$, and $\nu_{(p)}$ were optimized within realistic ranges: $2.00.01<\alpha_{(p)}<$ 2.0, $-0.99<\rho_{(p)}<0.0$, and $1.50.05<\nu_{(p)}<$ 1.5.
+In practice, fixing the backbone parameter $\beta$ is not restrictive for smile fitting.
+The remaining SABR parameters $(\alpha,\rho,\nu)$ provide sufficient flexibility to reproduce the observed smile with high accuracy (Zhang and Fabozzi [40]), while improving calibration stability under sparse data.
+We fixed $\beta=0.5$ throughout our experiments.
+Robustness with respect to the choice of $\beta$ is assessed in Section 5.4 (Figure 8).
+The parameters $\alpha_{(p)}$, $\rho_{(p)}$, and $\nu_{(p)}$ were optimized within realistic ranges: $0.01<\alpha_{(p)}<2.0$, $0.99<\rho_{(p)}<0.0$, and $0.05<\nu_{(p)}<1.5$.
 This yielded calibrated parameters $\{(\alpha_{(p)},\rho_{(p)},\nu_{(p)})\}_{p=1}^{P}$ for each observed maturity slice.
@@
   • If $\tilde{\tau}$ matches a calibrated maturity $\tau_{(p)}$, we used the calibrated parameters $(\alpha_{(p)},\rho_{(p)},\nu_{(p)})$.
-  • If $\tilde{\tau}$ is between two calibrated maturities, $\tau_{(p)}<\tilde{\tau}<\tau_{(p+1)}$, we used piecewise linear interpolation to get parameters $(\alpha_{\tilde{\tau}},\rho_{\tilde{\tau}},\nu_{\tilde{\tau}})$.
-For example, for $\alpha$:
+[…]
     $\alpha_{\tilde{\tau}}=\alpha_{(p)}+\frac{\tilde{\tau}-\tau_{(p)}}{\tau_{(p+1)}-\tau_{(p)}}(\alpha_{(p+1)}-\alpha_{(p)})$ (7)
@@
 In practice, Gaussian noise $\epsilon_{\mathcal{S}}\sim\mathcal{N}(0,\sigma^{2}_{\text{syn}})$ is added to the synthetic volatilities to prevent the model from becoming overly dependent on the synthetic data.
-3.3 Multi-Task Gaussian Process Formulation
-To combine information from the dense synthetic data (SABR-generated) and sparse market observations, we used a Multi-Task Gaussian Process (MTGP) framework.
-Standard GP models model a single output, while MTGP models handle multiple related outputs by capturing correlations between tasks [10, 11].
+3.3 Multitask Gaussian Process Formulation
+To combine information from the dense synthetic data and sparse market observations, we used a Multitask Gaussian process (MTGP) framework.
+A Gaussian process (GP) is a Bayesian nonparametric approach that can be used in regression problems [38].
+A GP defines a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$ such that any finite set of function values has a joint Gaussian distribution.
+Standard GP models handle a single output, while MTGP models handle multiple related outputs by capturing correlations between tasks [5, 2].
 We designated the synthetic SABR data $\mathcal{D}_{\mathcal{S}}$ as the source task and the sparse market data $\mathcal{D}_{\mathcal{T}}$ as the target task.
@@
 The parameters $\kappa_{\mathcal{S}}^{2}$ and $\kappa_{\mathcal{T}}^{2}$ control the amount of task-specific variance.
-For the shared components, we use learned task embeddings to infer the relationship. We modeled $h_{\mathcal{S}}$ and $h_{\mathcal{T}}$ as jointly Gaussian with a covariance structure:
+For the shared components, we used learned task embeddings to infer the relationship. We modeled $h_{\mathcal{S}}$ and $h_{\mathcal{T}}$ as jointly Gaussian with a covariance structure:
 $\text{Cov}(h_{\mathcal{Z}}(\mathbf{x}),h_{\mathcal{Z}^{\prime}}(\mathbf{x}^{\prime}))=\underbrace{\sigma^{2}_{h}\exp\left(-\frac{\|\mathbf{e}_{\mathcal{Z}}-\mathbf{e}_{\mathcal{Z}^{\prime}}\|^{2}}{l_{h}^{2}}\right)}_{\eqqcolon\,\hat{C}_{\mathcal{Z},\mathcal{Z}^{\prime}}}k(\mathbf{x},\mathbf{x}^{\prime})\quad\text{for}\quad\mathcal{Z,Z^{\prime}}\in\{\mathcal{S,T}\}$ (11)
-The term $\hat{C}_{\mathcal{Z},\mathcal{Z}^{\prime}}$ depends on task embeddings $\mathbf{e}_{\mathcal{Z}},\mathbf{e}_{\mathcal{Z}^{\prime}}\in\mathbb{R}^{d^{\prime}}$ ($d^{\prime}$ is a hyperparemeter that need to be predetermined) , a shared variance scaling parameter $\sigma^{2}_{h}$ and a length-scale parameter $l_{h}$.
+The term $\hat{C}_{\mathcal{Z},\mathcal{Z}^{\prime}}$ depends on task embeddings $\mathbf{e}_{\mathcal{Z}},\mathbf{e}_{\mathcal{Z}^{\prime}}\in\mathbb{R}^{d^{\prime}}$ ($d^{\prime}$ is a hyperparameter that needs to be predetermined), a shared variance scaling parameter $\sigma^{2}_{h}$ and a length-scale parameter $l_{h}$.
 This allowed the model to learn the level of information sharing.
@@
 $\mathbf{K}=\begin{bmatrix}k(\mathbf{X}_{\mathcal{S}},\mathbf{X}_{\mathcal{S}})&k(\mathbf{X}_{\mathcal{S}},\mathbf{X}_{\mathcal{T}})\\ k(\mathbf{X}_{\mathcal{T}},\mathbf{X}_{\mathcal{S}})&k(\mathbf{X}_{\mathcal{T}},\mathbf{X}_{\mathcal{T}})\end{bmatrix},\quad\mathbf{C}=\begin{bmatrix}C_{\mathcal{S,S}}\mathbf{1}_{M\times M}&C_{\mathcal{S,T}}\mathbf{1}_{M\times N}\\ C_{\mathcal{T,S}}\mathbf{1}_{N\times M}&C_{\mathcal{T,T}}\mathbf{1}_{N\times N}\end{bmatrix}$ (14)
-Where $k(\mathbf{X}_{\mathcal{Z}},\mathbf{X}_{\mathcal{Z^{\prime}}})$ is the matrix of kernel evaluations between input points, and $\mathbf{1}_{R\times H}$ is an $R\times H$ matrix of ones.
+where $k(\mathbf{X}_{\mathcal{Z}},\mathbf{X}_{\mathcal{Z^{\prime}}})$ is the covariance matrix, and $\mathbf{1}_{R\times H}$ is an $R\times H$ matrix of ones.
 3.4 Learning Task Relationships via Hierarchical Regularization
@@
 However, optimizing these embeddings can be challenging when directly maximizing the marginal likelihood, especially with imbalanced datasets.
-When the target dataset $\mathcal{D}_{\mathcal{T}}$ is much sparser than the source dataset $\mathcal{D}_{\mathcal{S}}$ (our typical scenario), standard maximum marginal likelihood estimation (MLE) can lead to two undesirable outcomes:
+When the target dataset $\mathcal{D}_{\mathcal{T}}$ is much sparser than the source dataset $\mathcal{D}_{\mathcal{S}}$, standard maximum marginal likelihood estimation (MLE) can lead to two undesirable outcomes:
   1. 1.
@@
 Instead of treating task embeddings as independent parameters to be learned solely through the likelihood, we assumed they are drawn from a common prior distribution, governed by hyperparameters $\bm{\mu}_{\text{e}}$ and $\sigma^{2}_{\text{e}}$:
-$P(\mathbf{e}_{\mathcal{Z}}|\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}})=\mathcal{N}(\mathbf{e}_{\mathcal{Z}}|\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}}\mathbf{I}_{d})\quad\text{for }\mathcal{Z}\in\{\mathcal{S},\mathcal{T}\}.$ (17)
-Here, $\bm{\mu}_{\text{e}}\in\mathbb{R}^{d^{\prime}}$ and $\sigma^{2}_{\text{e}}$ controlled the squared distance between task embeddings and their mean.
+[…]
+Here, $\bm{\mu}_{\text{e}}\in\mathbb{R}^{d^{\prime}}$ and $\sigma^{2}_{\text{e}}$ controls the squared distance between task embeddings and their mean.
 We learned both the embeddings $\mathbf{e}_{\mathcal{S}},\mathbf{e}_{\mathcal{T}}$ and the hyperparameters $\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}}$ from data.
@@
 $=$ $\sum_{\mathcal{Z}\in\{\mathcal{S},\mathcal{T}\}}\left(\frac{\|\mathbf{e}_{\mathcal{Z}}-\bm{\mu}_{\text{e}}\|^{2}}{2\sigma^{2}_{\text{e}}}+\frac{d^{\prime}}{2}\log\sigma^{2}_{\text{e}}+\frac{d^{\prime}}{2}\log(2\pi)\right)$ (17)
-The negative log marginal likelihood $-\log P(\mathbf{y}|\mathbf{X},\bm{\theta})$, derived from the joint Gaussian distribution of observations $\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})$ under the MTGP model, is given exactly by:
-$-\log P(\mathbf{y}|\mathbf{X},\bm{\theta})=\frac{1}{2}\mathbf{y}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{y}+\frac{1}{2}\log|\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma}|+\frac{M+N}{2}\log(2\pi)$ (20)
+The negative log marginal likelihood $\log P(\mathbf{y}|\mathbf{X},\bm{\theta})$, derived from the joint Gaussian distribution of observations $\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})$ under the MTGP model, is given exactly by:
+$\log P(\mathbf{y}|\mathbf{X},\bm{\theta})=\frac{1}{2}\mathbf{y}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{y}+\frac{1}{2}\log|\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma}|+\frac{M+N}{2}\log(2\pi)$ (18)
 where $\mathbf{K}\circ\mathbf{C}$ is the prior covariance matrix of the latent functions at training points ($M$ source, $N$ target), and $\bm{\Sigma}=\text{diag}(\sigma_{\mathcal{S}}^{2}\mathbf{1}_{M},\sigma_{\mathcal{T}}^{2}\mathbf{1}_{N})$ is the diagonal noise covariance matrix.
-Substituting the negative log marginal likelihood (NLML, Equation 20) and the relevant negative log prior (Equation 3.4) into the MAP objective (Equation 18), we arrived at the final objective function $\mathcal{L}(\bm{\theta})$ to minimize:
+Substituting the negative log marginal likelihood (NLML, equation (18)) and the relevant negative log prior (equation (3.4)) into the MAP objective (equation (16)), we arrived at the final objective function $\mathcal{L}(\bm{\theta})$ to minimize:
 $\mathcal{L}(\bm{\theta})=$ $\underbrace{\frac{1}{2}\mathbf{y}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{y}+\frac{1}{2}\log|\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma}|}_{\text{Data Fidelity Term (NLML)}}$
@@
 Log Prior)}}\underbrace{+\frac{M+N+2d^{\prime}}{2}\log(2\pi)}_{\text{Constant Term}}$ (19)
-We optimized Equation 21 using L-BFGS to find:
+We optimized equation (19) using L-BFGS to find:
   • Optimal task embeddings $\mathbf{e}_{\mathcal{S}}$ and $\mathbf{e}_{\mathcal{T}}$
@@
   • Other model hyperparameters (kernel parameters, noise variances, etc.)
-The optimization balanced fitting the data well (minimizing the NLML term) while keeping task embeddings plausible under their common prior (minimizing the regularization term).
+The optimization balanced fitting the data well while keeping task embeddings plausible under their common prior.
 The regularization strength adapted based on the learned variance $\sigma^{2}_{\text{e}}$.
@@
 Once we optimized all model parameters, we could make predictions for the target implied volatility surface at any new test point.
-The prediction followed standard GP principles, adapted for our multi-task framework.
+The prediction followed standard GP principles, adapted for our multitask framework.
 Given optimized parameters $\hat{\bm{\theta}}$ and the combined training dataset $\mathcal{D}=\mathcal{D}_{\mathcal{S}}\cup\mathcal{D}_{\mathcal{T}}$, we computed the predictive distribution for the target function $f_{\mathcal{T}}$ at a new test point $\mathbf{x}_{\ast}=[K_{\ast},\tau_{\ast}]^{\intercal}$.
 This involved the joint Gaussian distribution of the target function value $f_{\mathcal{T}}(\mathbf{x}_{\ast})$ and all training observations $\mathbf{y}$:
-$\begin{bmatrix}\mathbf{y}\\ f_{\mathcal{T}}(\mathbf{x}_{\ast})\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mathbf{0}_{M+N}\\ 0\end{bmatrix},\begin{bmatrix}\mathbf{K}\circ\mathbf{C}+\bm{\Sigma}&\mathbf{k}_{\ast}\\ \mathbf{k}_{\ast}^{\intercal}&C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\ast})\end{bmatrix}\right)$ (22)
+[…]
 The cross-covariance vector $\mathbf{k}_{\ast}$ related the test point to all training points across both tasks, incorporating learned task relationships via $C_{\mathcal{T,S}}$ and $C_{\mathcal{T,T}}$:
-$\mathbf{k}_{\ast}=\Big{[}\underbrace{C_{\mathcal{T,S}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{S},1}),\dots,C_{\mathcal{T,S}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{S},M})}_{\text{Covariance with source (SABR) points}},\underbrace{C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{T},1}),\dots,C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{T},N})}_{\text{Covariance with target (market) points}}\Big{]}^{\intercal}$ (23)
+[…]
 The factor $C_{\mathcal{T,S}}$ controlled the influence of synthetic SABR data, while $C_{\mathcal{T,T}}$ determined the weight of the market data.
@@
 The mean $\mu_{\ast}$ was our estimation of implied volatility at $\mathbf{x}^{\ast}$, and the variance $\sigma_{\ast}^{2}$ quantified prediction uncertainty.
-4 Experiments Design
+4 Experimental Design
 To evaluate SABR-MTGP against benchmark methods, we designed numerical experiments.
-Since the goal was to assess implied volatility surface construction accuracy, we needed a reliable ground truth reflecting realistic market behavior.
+Since the goal was to assess implied volatility surface construction accuracy, we required a reliable ground truth reflecting realistic market behavior.
 We chose the Heston stochastic volatility model [25] for this purpose, as it generates complex volatility dynamics similar to real markets.
-Our experimental setup included generating realistic ground truth data, creating sparse “market" observations, calibrating the SABR model, training the different models, and evaluating their performance across market conditions.
-This section details the methodology; Section 5 presents the findings.
+Our experimental setup comprised the generation of ground-truth data, construction of sparse “market” observations, slice-wise SABR calibration, model training, and evaluation across market regimes.
+This section details the methodology, and Section 5 presents the findings.
 4.1 Data Generation Framework
-Our experiments used two main datasets: (1) a sparse “market" dataset simulating real option price observations, and (2) a denser synthetic dataset derived from fitting a SABR model to these observations.
-The first was the target task, the second was the source task for multi-task learning.
+We used two datasets: a sparse “market” dataset and a denser synthetic dataset obtained by SABR calibration on the market dataset.
+The market dataset serves as the target task, and the synthetic dataset serves as the source task for multitask learning.
 4.1.1 Heston Ground Truth Data
-We generated synthetic “market" data using the Heston model [12], which describes asset price $S_{t}$ and variance $v_{t}$ dynamics via coupled SDEs:
+We generated synthetic “market” data using the Heston model [25], which describes asset price $S_{t}$ and variance $v_{t}$ dynamics via SDEs:
 $dS_{t}$ $=(r-q)S_{t}dt+\sqrt{v_{t}}S_{t}dW_{t}^{(1)}$ (25)
@@
 Brownian motions $W_{t}^{(1)}$ and $W_{t}^{(2)}$ have correlation $\rho$.
-We chose the Heston model because it generates realistic volatility smiles/skews and allows efficient option pricing via its characteristic function for accurate ground truth calculation.
+We chose Heston for its ability to reproduce realistic smiles and skews.
+Its characteristic function based pricing enables high-precision computations for ground-truth generation.
 Moreover, its parameters have financial interpretations, enabling simulation of various market regimes.
@@
 We set $S_{0}=100$, $r=0.03$, and $q=0.01$.
-To simulate “market" option data, we generated a dataset $\mathcal{D}_{\mathcal{T}}$ containing $N=166$ European call option contracts.
-We chose the maturities to match typical market distribution.
+To simulate “market” option data, we generated a dataset $\mathcal{D}_{\mathcal{T}}$ comprising $N=166$ European call option contracts.
+We designed maturities to mirror market availability, which is dense at short horizons and sparser at longer horizons.
 This led to expirations about monthly up to 9 months (0.08 to 0.75 years).
 Then, we used sparser intervals close to quarterly steps up to 1.75 years.
-And we included annual expirations at 2.0, 2.5, and 3.0 years.
+Additionally, we included annual expirations at 2.0, 2.5, and 3.0 years.
 For each maturity, we chose strike prices ($K$) to cover a specific moneyness range ($K/S_{0}$).
-This range depended on the maturity: $(0.7,1.6)$ for short-term ($\tau\leq 0.5$ years), and $(0.8,1.4)$ for mid-term ($0.5<\tau\leq 1.5$ years) and long-term ($\tau>1.5$ years).
+This range depended on the maturity: $[0.7,1.6]$ for short-term ($\tau\leq 0.5$ years), and $[0.8,1.4]$ for mid-term ($0.5<\tau\leq 1.5$ years) and long-term ($\tau>1.5$ years).
 Inside these ranges, we changed the strike intervals in a planned way.
-So, there were more strikes near the at-the-money (ATM) region (from $0.95S_{0}$ to $1.05S_{0}$) and fewer strikes further away.
+Consequently, there were more strikes near the at-the-money (ATM) region (from $0.95S_{0}$ to $1.05S_{0}$) and fewer strikes further away.
 Specifically, the strike intervals around ATM, near ATM, and far from the money were (2.5, 5.0, 15.0) for short-term, (5.0, 10.0, 25.0) for mid-term, and (10.0, 25.0, 50.0) for long-term maturities.
-Figure 1 showed the data point distribution for our experiments.
-Figure 1: Distribution of generated Heston “market" data points (target task $\mathcal{D}_{\mathcal{T}}$) in our experiments, by strike price and time to maturity.
+Figure 1 shows the data point distribution for our experiments.
+Figure 1: Distribution of generated Heston “market” data points (target task $\mathcal{D}_{\mathcal{T}}$) in our experiments, by strike price and time to maturity.
 Dashed red lines indicate evaluation maturities.
-For each contract, we computed the Heston price via Carr-Madan fast Fourier transform method [32].
-Then, we obtained the Black-Scholes implied volatility $\sigma_{\text{mkt}}(K,\tau)$ by inverting the Heston price using a root-finding algorithm.
-This produced the “market" dataset $\mathcal{D}_{\mathcal{T}}=\{(\mathbf{X}_{\mathcal{T}},\mathbf{y}_{\mathcal{T}})\}$, the target task data.
-4.1.2 SABR-Derived Synthetic Data
+For each option, we computed the Heston price via the Carr-Madan fast Fourier transform method (FFT) [6] using standard damping $\alpha=2.5$ and a high-resolution grid with $N=2^{21}$ and $c=12800$.
+We verified numerical stability by doubling $N$ and $c$, which changed call prices by less than $10^{-6}$.
+Here, $N$ denotes the number of frequency samples (FFT grid size) and $c$ is the integration-range truncation parameter in Carr-Madan FFT.
+The spacing $\eta=c/N$ and the log-strike span $\Lambda=\tfrac{2\pi}{N\eta}$ in the Carr-Madan formulation control the tradeoff between aliasing and accuracy.
+IVs are computed via a safeguarded Newton method (tolerance $10^{-5}$, maximum 200 iterations, $\sigma\in[10^{-4},5]$).
+This produced the “market” dataset $\mathcal{D}_{\mathcal{T}}=\{(\mathbf{X}_{\mathcal{T}},\mathbf{y}_{\mathcal{T}})\}$, the target task data.
+4.1.2 SABR-generated Synthetic Data
 The SABR-MTGP used a denser synthetic dataset derived by fitting the SABR model to the Heston data.
-This involved several steps.
-First, for each unique maturity $\tau_{(p)}\in\mathbb{T}_{\text{mkt}}$, SABR parameters ($\alpha_{(p)},\rho_{(p)},\nu_{(p)}$) were calibrated by minimizing MSE between SABR volatilities $\sigma_{\text{SABR}}$ and market volatilities $\sigma_{\text{mkt}}$ from $\mathcal{D}_{\mathcal{T}}$ ($\beta=0.5$ fixed).
-We used Nelder-Mead optimization with constraints, yielding calibrated parameters $\{(\alpha_{(p)},\rho_{(p)},\nu_{(p)})\}$ for the set $\mathbb{T}_{\text{mkt}}$.
-Second, we obtained SABR parameters for any time-to-maturity $\tilde{\tau}$ via piecewise linear interpolation (Equation 9) between parameters at the nearest surrounding calibrated maturities in $\mathbb{T}_{\text{mkt}}$.
-Constant extrapolation was used outside this range.
+We calibrated the SABR following the scheme in Section 3.2.
 Figure 2 shows the calibrated SABR parameters across maturities for the Base Heston scenario.
-The parameters showed term structures consistent with market behavior, supporting the interpolation.
 Figure 2: Evolution of calibrated SABR parameters ($\alpha,\beta,\rho,\nu$) across maturity slices for the Base Heston scenario. Note $\beta$ was fixed at 0.5 during calibration.
-Third, we defined a dense $35\times 35$ grid $(\tilde{K}_{j},\tilde{\tau}_{j})$, covering strike range ($K\in[70,160]$) and maturity range ( $\tau\in[0.08,3.0]$).
-At each grid point, we used the interpolated/extrapolated SABR parameters with the Hagan formula (Equation 7) to compute synthetic implied volatility $y_{\mathcal{S},j}=\sigma_{\text{SABR}}(\tilde{K}_{j},F_{\tilde{\tau}_{j}},\tilde{\tau}_{j};\alpha_{\tilde{\tau}_{j}},\beta,\rho_{{\tilde{\tau}}_{j}},\nu_{{\tilde{\tau}}_{j}}).$
+Then, we defined a dense $35\times 35$ grid $(\tilde{K}_{j},\tilde{\tau}_{j})$, covering strike range ($K\in[70,160]$) and maturity range ($\tau\in[0.08,3.0]$).
+At each grid point, we used the Hagan formula equation (5) to compute synthetic implied volatility $y_{\mathcal{S},j}=\sigma_{\text{SABR}}(\tilde{K}_{j},F_{\tilde{\tau}_{j}},\tilde{\tau}_{j}).$
 Finally, we added small Gaussian noise $\epsilon_{\mathcal{S}}\sim\mathcal{N}(0,0.01^{2})$ to the synthetic data.
@@
 4.2 Comparative Models and Training
-We compared three approaches for constructing the IVS.
-First, SABR Interpolation, a parametric baseline using calibrated and interpolated SABR parameters with the Hagan formula (Equation 7).
-Second, a Single-Task Gaussian Process (GP) trained only on the sparse Heston “market" data $\mathcal{D}_{\mathcal{T}}$.
-This GP used a Matérn 5/2 kernel on $(K,\tau)$ inputs and learned hyperparameters via L-BFGS optimization of marginal likelihood.
-Third, the proposed SABR-Informed Multi-Task GP (SABR-MTGP), our transfer learning framework.
-It was trained jointly on synthetic source data $\mathcal{D}_{\mathcal{S}}$ and market target data $\mathcal{D}_{\mathcal{T}}$, using the ICM structure with a shared Matérn 5/2 input kernel and a task kernel based on learned one-dimensional embeddings.
+We evaluated six methods: (i) SABR [^1] Hagan asymptotics with fixed $\beta=0.5$; slice-wise calibration of $(\alpha,\rho,\nu)$., (ii) Single-Task GP [^2] Matérn 5/2 kernel on $(K,\tau)$ with L-BFGS training, (iii) the proposed SABR-Informed Multitask GP (SABR-MTGP) [^3] ICM structure with a shared Matérn 5/2 input kernel and a task kernel based on learned one-dimensional embeddings; optimized by L-BFGS., (iv) arbitrage-free SSVI [17], (v) Cubic Spline, and (vi) Neural Network with SABR pre-training [^4] Two-hidden-layer MLP (64 units each, ReLU) pre-trained on synthetic SABR data and fine-tuned on target data with Adam optimizer..
 4.3 Evaluation Framework and Robustness Analysis
-We evaluated model performance against the Heston ground truth on a predefined evaluation grid for fair comparison.
+We evaluated model performance against the Heston ground truth on a predefined evaluation grid for a fair comparison.
 This grid, separate from training data, included points $(K_{\ast},\tau_{\ast})$ spanning a relevant moneyness range (e.g., $K_{\ast}/S_{0}\in[0.8,1.4]$) at specific test maturities $\tau_{\ast}$ (e.g., 0.3, 0.9, 2.2 years) representing near-, mid-, and long-term.
-For each evaluation point, we computed the ground truth Heston IV using a high-precision FFT approach with the Heston characteristic function, followed by numerical inversion to obtain corresponding implied volatility.
-Each model (SABR Interpolation, GP, SABR-MTGP) predicted $\hat{\sigma}(K_{\ast},\tau_{\ast})$ on this grid.
-We measured performance using Root Mean Squared Error (RMSE) and Mean Absolute Error (MeanAE) against the ground truth $\sigma_{\text{Heston}}(K_{\ast},\tau_{\ast})$, calculated over valid predictions.
+For each point, we computed Heston prices via the Carr-Madan FFT and inverted to IVs by safeguarded Newton.
+Each model (SABR, GP, SABR-MTGP, SSVI, Cubic Spline, and NN) predicted $\hat{\sigma}(K_{\ast},\tau_{\ast})$ on this grid.
+We measured performance using Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) against the ground truth $\sigma_{\text{Heston}}(K_{\ast},\tau_{\ast})$, calculated over valid predictions.
 To assess robustness, we repeated experiments using ten Heston parameter configurations (Table 1), including the Base scenario.
 These settings simulated various market dynamics by altering parameters like mean reversion $\kappa$, volatility levels $\theta,\nu_{\text{vol}}$, correlation $\rho$, and initial term structure $v_{0}$ vs $\theta$.
-This allowed analysis across different market conditions.
-Robustness evaluation focused on metrics at distinct test maturities.
-We also analyzed the MTGP’s learned task correlation and variance decomposition for each scenario to understand its adaptive information-sharing.
 Table 1: Heston Parameter Configurations for Robustness Analysis. Common parameters are $S_{0}=100,r=0.03,q=0.01$.
@@
 5 Results and Discussion
-This section presents our findings on SABR-MTGP performance compared with SABR interpolation and the standard GP.
-We first analyze performance under the Base Heston scenario, then examine robustness across various market conditions.
-5.1 Performance in the Base Scenario
-Figures 3(a), 3(b), and 3(c) show fitted implied volatility curves from the three models compared against Heston ground truth, for near-term ($\tau=0.30$), mid-term ($\tau=0.90$), and long-term ($\tau=2.20$) maturities in the Base scenario.
-Each figure includes fitted curves and residuals.
-(a) Near-term ($\tau=0.30$).
-(b) Mid-term ($\tau=0.90$).
-(c) Long-term ($\tau=2.20$).
-Figure 3: Implied volatility curve comparison and residuals under the Base Heston scenario for (a) Near-term, (b) Mid-term, and (c) Long-term maturities.
-The results showed different performance patterns across the term structure.
-For near-term maturity (Figure 3(a)), where market data is relatively dense, SABR-MTGP achieved the most accurate fit across strikes (RMSE=0.0014).
-The standard GP (RMSE=0.0024) showed systematic deviations, suggesting difficulty capturing the precise smile pattern.
-Direct SABR interpolation (RMSE=0.0057) struggled with curvature, showing systematic errors consistent with limitations of its parametric form in matching these Heston dynamics.
-Moving to mid-term maturity (Figure 3(b)), data sparsity increased.
-SABR-MTGP continued to provide the closest fit (RMSE=0.0005).
-SABR interpolation also performed very well (RMSE=0.0009), capturing the overall shape effectively.
-However, the standard GP (RMSE=0.0043) showed larger and more scattered errors, particularly for strikes away from the money.
-For long-term maturity (Figure 3(c)), SABR interpolation performed best according to RMSE (0.0005), benefiting from its structure to provide a reasonable approximation.
-SABR-MTGP also achieved very good accuracy (RMSE=0.0007).
-The standard GP (RMSE=0.0025) exhibited notably higher errors compared to the other two methods, with residuals spanning approximately from $-0.006$ to $+0.002$ volatility points, indicating difficulties in generalizing from the sparse long-term data.
-These findings highlight the advantages of our hybrid approach.
-In data-rich regions (near-term), SABR-MTGP combined market data with structural guidance to outperform alternatives.
-As data gets sparser (mid- to long-term), it used SABR structure effectively while retaining flexibility to adapt to Heston dynamics, balancing SABR’s rigidity and standard GP’s data dependence.
-5.2 Robustness Across Market Conditions
-The performance heatmaps in Figures 4, 5, and 6 summarized model accuracy (RMSE and MeanAE) across the ten Heston parameter settings, evaluated at near-, mid-, and long-term horizons.
-Lighter shades indicated lower errors.
-Figure 4: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Near-term maturity ($\tau=0.3$).
-Figure 5: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Mid-term maturity ($\tau=0.9$).
-Figure 6: Performance heatmap (RMSE and MeanAE) across different Heston parameter settings for the Long-term maturity ($\tau=2.2$).
-These heatmaps show consistent performance patterns.
-SABR-MTGP consistently demonstrated strong performance, frequently ranking as the best or a very close second-best method.
-The standard GP showed reasonable performance in the near-term, but its accuracy dropped significantly at long-term maturities, where it yielded the highest errors.
-Conversely, SABR interpolation struggled in the near-term, often producing the highest errors, but benefited from its fixed structure in long-term cases.
-It frequently achieved the lowest errors, especially with very sparse data.
-SABR-MTGP’s ability to adaptively blend structural guidance with data-driven flexibility allowed it to maintain high accuracy overall, overcoming SABR’s near-term issues and the GP’s long-term instability.
+This section reports cross regime robustness and then uses Base scenario slices as diagnostics. We compare SABR, the single-task GP, SABR-MTGP, SSVI, cubic spline, and a neural network baseline.
+5.1 Performance and Robustness Across Market Conditions
+Tables 2 and 3 summarize model accuracy (RMSE and MAE) across ten Heston settings at near-term and long-term horizons.
+Values are scaled by $10^{-3}$, and the best metric value per setting is shown in bold.
+At the short maturity, SABR-MTGP is best in eight settings and GP leads in two.
+At the long maturity, the leader alternates by regime, with SABR best in six settings, SABR-MTGP best in three, and GP best in one.
+SSVI and cubic spline are competitive in a few settings yet they are not consistently dominant.
+The neural network baseline is generally weaker.
+Table 2: Error summary at Near-term maturity ($\tau=0.3$). Values are scaled ($\times 10^{-3}$).
+Setting SABR GP SABR-MTGP SSVI Cubic Spline NN
+RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
+Base 5.78 5.22 1.93 1.63 1.69 1.44 4.27 3.42 2.11 1.80 4.67 3.49
+Moderate Mean-Rev 4.50 4.04 1.55 1.29 1.42 1.25 4.09 3.17 1.79 1.54 6.31 4.36
+Low Mean-Rev 6.60 5.97 2.24 1.79 2.09 1.83 4.34 3.39 2.28 1.92 4.91 3.75
+High Vol-Regime 4.99 4.56 0.83 0.69 0.71 0.61 3.17 2.50 1.44 1.10 2.67 1.98
+Low Vol-Regime 6.74 6.36 4.15 2.48 3.94 2.35 4.04 3.15 5.04 3.17 10.89 7.37
+Moderate Correlation 4.10 3.72 1.38 1.08 1.23 1.05 1.90 1.71 1.93 1.45 3.21 2.66
+Strong Correlation 7.26 6.13 4.28 3.25 3.10 2.41 7.04 5.44 3.84 3.03 11.51 7.41
+Term Structure Up 4.88 4.37 1.34 1.14 1.59 1.41 4.42 3.34 1.84 1.57 5.02 3.60
+Term Structure Down 5.24 4.76 1.12 0.96 0.82 0.72 6.25 4.75 1.43 1.09 3.57 2.43
+Mixed Regime 2.16 2.00 0.38 0.29 0.89 0.65 1.12 0.83 0.75 0.56 2.03 1.52
+Table 3: Error summary at Long-term maturity ($\tau=2.2$). Values are scaled ($\times 10^{-3}$).
+Setting SABR GP SABR-MTGP SSVI Cubic Spline NN
+RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
+Base 0.42 0.27 0.82 0.69 0.61 0.59 0.45 0.33 1.08 0.66 4.60 4.10
+Moderate Mean-Rev 0.36 0.26 1.18 0.99 0.86 0.72 0.83 0.73 0.39 0.27 3.44 2.87
+Low Mean-Rev 0.55 0.45 1.60 1.17 1.81 1.56 1.60 1.32 3.15 1.96 5.24 4.48
+High Vol-Regime 0.33 0.24 0.24 0.19 0.17 0.14 0.90 0.76 0.43 0.31 4.05 3.53
+Low Vol-Regime 0.51 0.32 2.52 1.73 0.38 0.33 0.75 0.59 1.17 0.72 3.66 2.90
+Moderate Correlation 0.12 0.10 1.24 1.19 0.35 0.27 0.65 0.59 1.18 0.79 4.99 4.30
+Strong Correlation 0.74 0.41 1.36 0.67 0.15 0.12 0.83 0.74 0.58 0.34 4.01 3.54
+Term Structure Up 0.42 0.29 0.32 0.20 0.95 0.83 0.49 0.40 0.49 0.32 3.92 3.25
+Term Structure Down 0.57 0.54 0.97 0.84 0.68 0.47 2.34 1.87 1.69 1.22 3.66 3.24
+Mixed Regime 0.10 0.07 0.49 0.46 0.13 0.11 0.40 0.34 0.34 0.25 4.53 3.96
+Figures 3 and 4 provide visual evidence from the Base scenario.
+In the near-term slice, SABR shows a systematic error pattern relative to the Heston truth that is consistent with the limitations of a rigid parametric form.
+The single-task GP follows the curve yet exhibits small local oscillations in the wings where observations are sparse.
+SABR-MTGP tracks the truth closely and its residuals are centered and smooth.
+Figure 3: Implied volatility curve comparison and residuals under the Base Heston scenario for near-term
+Figure 4: Implied volatility curve comparison and residuals under the Base Heston scenario for long-term
+In the long-term slice, the three methods are closer. SABR is often the best while SABR-MTGP remains very close and smooth. The single-task GP shows mild undulations yet the deviations are small.
+These patterns align with the design of SABR-MTGP.
+Short maturities contain more observations and the SABR-based source task acts as a stabilizer that suppresses small oscillations of a pure GP while preserving flexibility.
+At long maturities data are sparse and a simple parametric structure can be competitive.
+SABR-MTGP uses learned task embeddings with hierarchical regularization to adapt the level of transfer.
+When SABR and the target regime align, the learned correlation increases (cf.
+Section 5.3).
+5.2 No-Arbitrage Analysis
+We evaluate no-arbitrage by computing the butterfly spread and total variance. For each maturity $\tau$ and ordered strikes $\{K_{i}\}_{i=1}^{100}\subset[80,140]$, we define the butterfly spread
+$B_{i}(\tau)\,=\,C(K_{i+1},\tau)\,-\,2\,C(K_{i},\tau)\,+\,C(K_{i-1},\tau),$ (27)
+which must satisfy $B_{i}(\tau)\geq 0$ (convexity in $K$).
+We report the butterfly violation rate as $\{i:B_{i}(\tau)<-\varepsilon\}/\#\{i\}$ with tolerance $\varepsilon=2\times 10^{-5}$.
+We define the total variance $w(k,\tau)=\sigma^{2}(k,\tau)\,\tau$ with forward $F_{\tau}$ and log-moneyness $k=\ln(K/F_{\tau})$.
+The total variance should be non-decreasing in $\tau$ at fixed $k$ if the IVS is calendar arbitrage-free.
+Tables 45 report violation rates, and Figure 5 visualizes butterfly spreads and total variance.
+Across Tables 45, SABR, SSVI, cubic spline, GP, and SABR-MTGP rarely exhibit butterfly violations on the evaluation grids; the few exceptions are small ($\leq 3.33\%$) and arise mainly for GP at long-term maturity and for cubic spline at near-term maturity.
+The NN baseline shows higher violation rates in several regimes, especially at long-term maturities.
+Figure 5 corroborates these findings: butterfly spreads of the proposed method are nonnegative, and total variance panels exhibit the expected increase in $w$ with maturity at fixed $k$ consistent with the absence of calendar arbitrage.
+Table 4: Near-term butterfly violation rate ($\tau=0.3$). $\times$ denotes no violation.
+Setting SABR SSVI Cubic Spline GP SABR-MTGP NN
+Base $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Moderate Mean-Rev $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Low Mean-Rev $\times$ $\times$ 1.11% $\times$ $\times$ $\times$
+High Vol-Regime $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Low Vol-Regime $\times$ $\times$ $\times$ $\times$ $\times$ 4.44%
+Moderate Correlation $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Strong Correlation $\times$ $\times$ 2.22% $\times$ $\times$ 1.11%
+Term Structure Up $\times$ $\times$ $\times$ $\times$ $\times$ 13.33%
+Term Structure Down $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Mixed Regime $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Table 5: Long-term butterfly violation rate ($\tau=2.2$). $\times$ denotes no violation.
+Setting SABR SSVI Cubic Spline GP SABR-MTGP NN
+Base $\times$ $\times$ $\times$ 3.33% $\times$ 25.56%
+Moderate Mean-Rev $\times$ $\times$ $\times$ $\times$ $\times$ 26.67%
+Low Mean-Rev $\times$ $\times$ $\times$ 1.11% $\times$ 22.22%
+High Vol-Regime $\times$ $\times$ $\times$ $\times$ $\times$ 37.78%
+Low Vol-Regime $\times$ $\times$ $\times$ $\times$ $\times$ 14.44%
+Moderate Correlation $\times$ $\times$ $\times$ $\times$ $\times$ 33.33%
+Strong Correlation $\times$ $\times$ $\times$ $\times$ $\times$ 24.44%
+Term Structure Up $\times$ $\times$ $\times$ $\times$ $\times$ 33.33%
+Term Structure Down $\times$ $\times$ $\times$ 1.11% $\times$ 21.11%
+Mixed Regime $\times$ $\times$ $\times$ $\times$ $\times$ 35.56%
+(a) Butterfly spreads across strikes for three maturities (Base).
+(b) Total variance $w(k,\tau)$ across methods (six panels, Base).
+Figure 5: No-arbitrage test.
+Panel (a) visualizes butterfly spreads; most methods remain above the zero line within a small tolerance band, while NN and GP occasionally dip negative at mid/long maturities.
+Panel (b) shows six panels of total variance $w(k,\tau)$ curves for SABR, SSVI, Cubic Spline, GP, SABR-MTGP, and NN; for most methods the level of $w$ increases with maturity at fixed $k$ and the $k$-profiles are smooth, with minor departures mainly for $N$ and GP.
 5.3 Learned Task Relationship
-The MTGP framework learned the relationship between the SABR source task and the Heston market target task.
-Figure 7 shows this adaptive relationship for three Heston settings, illustrating variations in learned correlation, covariance structure, and variance decomposition (task-specific proportion $\kappa^{2}_{\mathcal{Z}}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$ vs.
-shared variance proportion $\sigma^{2}_{h}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$).
-(a) Low Correlation Example: Moderate Mean-Rev Setting (Learned Correlation = 0.22).
-(b) Moderate Correlation Example: Base Setting (Learned Correlation = 0.41).
-(c) High Correlation Example: Term Structure Up Setting (Learned Correlation = 0.90).
-Figure 7: Examples of learned task relationship analysis for the SABR-MTGP model under different Heston parameter settings: (a) Moderate Mean-Rev, (b) Base, and (c) Term Structure Up.
-Panels show variations in learned correlation (left), covariance structure (center), and variance decomposition (right).
-The learned correlation varied with Heston parameters.
-Faster mean reversion (Moderate Mean-Rev) yielded low correlation (0.22), perhaps because SABR captured these dynamics less well.
-An upward term structure (Term Structure Up) resulted in high correlation (0.90), suggesting SABR provided a better structural approximation in this case.
-The Base setting showed moderate correlation (0.41).
-This showed the model’s ability to assess similarity between SABR information and target data based on market conditions.
-The task covariance matrices reflected these correlations.
-A key pattern was that the shared covariance component consistently accounted for most variance in both tasks across scenarios.
-Even if SABR is not a perfect match (low correlation), the MTGP leveraged shared properties like smoothness, preventing overfitting to sparse target data and performing better than the standard GP at longer maturities.
-When SABR aligns well (high correlation), the model strongly used this information, achieving accuracy competitive with or better than SABR interpolation, while maintaining flexibility.
-5.4 Parameter Sensitivity Analysis
-We further examined robustness by analyzing performance sensitivity to Heston parameter changes.
-Using the coefficient of variation (CV) of RMSE, Figure 8 showed how much the precision of each model fluctuated within the parameter categories relative to its average performance.
-Lower CV (lighter colors) indicated more stable performance.
-(a) Near-term ($\tau=0.3$).
-(b) Mid-term ($\tau=0.9$).
-(c) Long-term ($\tau=2.2$).
-Figure 8: Parameter sensitivity heatmaps showing the Coefficient of Variation (CV) of RMSE for GP, SABR, and SABR-MTGP across different Heston parameter categories (y-axis) for near-, mid-, and long-term evaluation maturities.
-Lower CV (lighter color) indicated greater performance stability within that parameter category.
-The sensitivity heatmaps highlighted model stability differences.
-The standard GP often showed high CV, indicating its performance was sensitive to underlying volatility dynamics, especially concerning ‘Correlation’ and ‘Vol-of-Vol’ changes at longer maturities.
-SABR interpolation typically had the lowest CV due to its fixed parametric form, but this stability could come at the cost of lower accuracy, particularly near-term.
-SABR-MTGP generally displayed intermediate stability.
-It demonstrated improved stability over the standard GP in the mid- and long-term.
-Although it was less stable than SABR interpolation.
-The sensitivity results thus complemented accuracy findings: SABR-MTGP found a middle ground, using the prior for improved stability over GP, while retaining adaptive flexibility unlike SABR interpolation.
+The MTGP framework learns the cross-task relationship between the SABR source and the Heston target.
+Figure 6 illustrates the learned task correlation, the task covariance, and the variance decomposition (task-specific proportion $\kappa^{2}_{\mathcal{Z}}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$ versus shared proportion $\sigma^{2}_{h}/(\sigma^{2}_{h}+\kappa^{2}_{\mathcal{Z}})$) of two representative settings.
+(a) Term Structure Down Setting.
+(b) Mixed Regime setting.
+Figure 6: Learned task relationship for two representative Heston settings: (a) Term Structure Down and (b) Mixed Regime. Each panel shows the task correlation, covariance, and variance decomposition.
+In the Term Structure Down case, the cross-task correlation is higher ($\approx 0.71$), indicating stronger transfer when the regime aligns with SABR’s structure.
+In the Mixed Regime, the lower correlation ($\approx 0.40$) implies greater reliance on task-specific variation.
+This adaptive behavior limits negative transfer while preserving the benefits of information transfer from SABR.
+5.4 Sensitivity Analyses
+5.4.1 Noise in the source task
+We study how noise $\epsilon_{\mathcal{S}}\sim\mathcal{N}(0,\sigma_{\text{syn}}^{2})$ in the SABR-generated source task affects model accuracy.
+We inject zero-mean Gaussian noise with standard deviation $\sigma_{\text{syn}}\in\{0,\,0.005,\,0.01,\,0.015,\,0.02\}$ into the synthetic source implied volatilities, and evaluate overall accuracy on the same out-of-sample grid used elsewhere in this section.
+Figure 7 reports the Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) aggregated over the evaluation grid for three methods: SABR-MTGP, single-task GP, and a neural network pretrained on the source and finetuned on the target (NN).
+Empirically, SABR-MTGP exhibits a clear U-shaped response: small amounts of noise reduce error and the best performance is attained around $0.01$, after which errors slowly increase.
+This behavior is consistent with the role of mild perturbations as a regularizer that reduces over-reliance on the synthetic dataset, while excessive noise reduces useful structural information.
+The NN baseline degrades monotonically with increasing $\sigma_{\text{syn}}$, reflecting its higher sensitivity to the quality of the pretraining dataset.
+Practically, these results suggest that modest source-task noise (here, $\sigma_{\text{syn}}$ in the range [0.005, 0.015][0.005,\,0.015]) can improve the robustness and generalization of SABR-MTGP.
+Figure 7: Sensitivity of overall accuracy to Gaussian noise injected into the SABR source labels.
+Left: RMSE vs.
+noise level $\sigma_{\text{syn}}$.
+Right: MAE vs.
+$\sigma_{\text{syn}}$.
+SABR-MTGP benefits from modest noise (minimum near $\sigma_{\text{syn}}\!\approx\!0.01$), and NN degrades as $\sigma_{\text{syn}}$ increases.
+5.4.2 Backbone parameter $\beta$
+In the previous sections, we treat the SABR backbone $\beta$ as a hyperparameter for generating synthetic data.
+In this section, we verify that our conclusions do not hinge on its value.
+We repeat the Base experiment for $\beta\in\{0,\,0.25,\,0.5,\,0.75,\,1.0\}$, and compute out-of-sample errors on the same evaluation grid.
+Figure 8 shows that SABR-MTGP is insensitive to $\beta$ within a broad range, and outperforms GP in all cases.
+The neural network baseline is less accurate.
+Figure 8: Effect of the SABR backbone parameter $\beta$ on overall accuracy in the Base setting. Left: RMSE vs. $\beta$. Right: MAE vs. $\beta$.
 5.5 Summary
-The experiments using Heston data suggested SABR-MTGP effectively constructed the IVS.
-By combining SABR’s structural information with market observations in a multi-task framework, it achieved better accuracy and stability than using either a standard GP or SABR.
-It effectively integrated structural information with observed data, adaptively leveraging SABR insights by assessing their relevance to the target dataset.
+Experiments on Heston data show that SABR-MTGP effectively constructs the IVS.
+By combining SABR’s structural information with market observations in a multitask framework, it achieves better accuracy and stability than either GP or SABR alone, adaptively leveraging SABR insights according to their relevance to the target data.
 6 Application to SPX Market Data
-To further evaluate the practical applicability of SABR-MTGP, we tested its performance on real-world market data for SPX options (provided by the OptionMetrics database), observed on August 1, 2023.
+To assess the practical applicability of SABR-MTGP, we evaluate it on SPX option data from the OptionMetrics database dated August 1, 2023.
 This case study complements the controlled experiments using the Heston model by assessing the method’s behavior under actual market conditions.
 6.1 Data
-The dataset included European call options on the SPX index.
-We preprocessed the data by filtering for valid quotes, calculating time-to-maturity, and finding implied volatilities from mid-prices.
-We used the Zero Coupon Bond yield curve as the interest rate curve, and estimated the dividend yield using put-call parity.
-This gave us several hundred valid market implied volatility observations across different strikes and maturities up to about one year.
+The dataset includes European call options on the SPX index.
+We preprocess the data by filtering for valid quotes, computing time-to-maturity, and computing implied volatilities from mid-prices.
+We use the zero-coupon yield curve for discounting and estimate the dividend yield via put-call parity.
+This gives us several hundred valid market implied volatility observations across different strikes and maturities up to about one year.
 Key statistics summarizing this dataset are in Table 6.
-This processed market data was the target task ($\mathcal{D}_{\mathcal{T}}$) for both the SABR-MTGP model.
+This processed market dataset serves as the target task ($\mathcal{D}_{\mathcal{T}}$) for all methods.
 Table 6: Descriptive Statistics for the Preprocessed SPX Call Option Dataset (August 1, 2023)
@@
 6.2 Results
-We do not have a ground truth IVS for real market data.
-So, the evaluation used a qualitative assessment of the model fits and the resulting surfaces.
 Figure 9 shows typical implied volatility slices for near-term ($\tau=0.2190$ years) and mid-term ($\tau=0.8898$ years) maturities.
@@
 (b) Mid-term maturity ($\tau=0.8898$ years).
-Figure 9: Implied volatility slices for SPX options on August 1, 2023.
-Comparison of SABR-MTGP, GP, and SABR interpolation against market data points.
-Residuals (Model - Market) are shown below each slice.
-In both slices (Figure 9), all three methods generally captured the observed smile/skew pattern of the market.
-But the residual plots showed important differences in fit quality.
-For the near-term slice (Figure 9(a)), SABR interpolation had systematic errors.
-It underestimated volatility for low strikes (ITM) and overestimated near the money.
-This was shown by the consistent pattern in its residuals.
-Both the standard GP and SABR-MTGP fitted the observed market points better, especially around the at-the-money region.
-But there was a notable difference in the deep in-the-money region (strikes 3250-4000).
-The standard GP curve, although sloping downward, showed a shape considered less regular and potentially less financially realistic compared to volatility curve given by the SABR.
-The SABR-MTGP curve maintained a smoother, decreasing volatility in this data-sparse region.
-This produced a more financially realistic interpolation, likely because of the structural guidance from the SABR source task.
-For the mid-term slice (Figure 9(b)), where the smile was flatter and the market data was denser, all models achieved a very good fit.
-Notably, the SABR model fit improved significantly compared to the near-term slice.
-The residuals for both GP and SABR-MTGP were particularly small (generally within +/- 0.001), demonstrating excellent agreement with market observations.
+Figure 9: Implied volatility slices for SPX options on August 1, 2023. Comparison of SABR-MTGP, GP, and SABR against market data points. Residuals (Model - Market) are shown below each slice.
+In both slices (Figure 9), all three methods generally capture the observed smile/skew pattern of the market.
+However, the residual plots reveal notable differences in fit quality.
+For the near-term slice (Figure 9(a)), SABR has systematic errors shown by the consistent pattern in its residuals.
+Both the GP and SABR-MTGP fit the observed market points better, especially around the at-the-money region.
+However, there is a notable difference in the deep in-the-money region (strikes 3250-4000).
+Although downward-sloping, the standard GP curve exhibits local irregularities that render it less financially plausible than the SABR-implied curve.
+The SABR-MTGP curve maintains a smoother, decreasing volatility in this data-sparse region.
+This produces a more financially realistic interpolation, likely because of the structural guidance from the SABR source task.
+For the mid-term slice (Figure 9(b)), where the smile is flatter and the market data is denser, all models achieve a very good fit.
+Notably, the SABR model fit improves significantly compared to the near-term slice.
+The residuals for both GP and SABR-MTGP are particularly small (generally within +/- 0.001), demonstrating excellent agreement with market observations.
 (a) GP
-(b) SABR Interpolation
+(b) SABR
 (c) SABR-MTGP
-Figure 10: SPX Implied Volatility Surfaces constructed using (a) GP, (b) SABR Interpolation, and (c) SABR-MTGP, based on market data from August 1, 2023.
-Black dots represent observed market implied volatilities.
-The three plots (Figure 10) showed the different shapes of the implied volatility surfaces produced by the three models.
-The standard GP surface (Figure 10(a)) matched the observed market data points very closely.
-However, in data-sparse regions, particularly for deep out-the-money options ($K$ is large), it showed some unusual patterns that were inconsistent with typical market behavior, such as a downward slope in volatility where an upward sloping skew is generally expected.
-The SABR surface (Figure 10(b)) was naturally smooth because it was based on a specific mathematical formula.
-But, because of this fixed structure, it might not have captured all the specific details or patterns seen in the market data.
-The SABR-MTGP surface (Figure 10(c)) seemed to offer a good balance between the other two methods.
-It fitted the market data well, similar to the standard GP.
-And, it maintained a smoother and more economically consistent shape compared to the standard GP, especially when extending to areas with less data, such as longer times-to-maturity or strike prices far from the current price.
-This visual comparison suggested that the SABR-MTGP model successfully used the general structure of the SABR model to guide the construction of the surface, particularly where market data was sparse.
-This helped to create a more stable and realistic surface overall.
-In summary, the application to real SPX market data showed the model’s ability to integrate structural financial knowledge with sparse empirical observations.
-This led to well-behaved and plausible implied volatility surfaces suitable for practical applications.
+(d) SSVI
+(e) Cubic Spline
+(f) NN
+Figure 10: SPX implied volatility surfaces constructed using six methods: GP, SABR, SABR-MTGP, SSVI, cubic spline, and a neural network. Black dots denote market observations.
+Across the six surfaces in Figure 10, SABR-MTGP produces a smooth, well-behaved surface that tracks market levels across maturities and remains stable in sparse regions.
+The single-task GP is competitive overall but exhibits mild local undulations near the wings and toward the long end.
+The SABR and the neural network are smooth but can be too rigid around the wings.
+SSVI delivers a shape-constrained surface, whereas the cubic spline is generally smooth but shows occasional edge bending and a localized bulge at mid-maturities and ATM.
+Overall, the figures illustrate the intended trade-off: SABR-MTGP balances structural guidance and flexibility, preserving smoothness and plausibility while remaining close to the market points.
+Figure 11: Butterfly spreads across strikes.
+The butterfly spreads shown in Figure 11 remain nonnegative for SABR-MTGP, SABR, SSVI and cubic spline on the inspected slices, while GP and NN exhibit occasional dips near mid or long maturities.
+These diagnostics indicate that the surface generated by the proposed method is consistent with static no-arbitrage on the SPX sample.
 7 Conclusion
 In this paper, we proposed SABR-MTGP, an approach bridging structural and data-driven methods for IVS construction.
-The key idea is framing IVS construction as a multi-task learning problem where we used structural information from the SABR model to improve predictions from sparse market data.
+The key idea is framing IVS construction as a multitask learning problem where we used structural information from the SABR model to improve predictions from sparse market data.
 Our approach offered several contributions.
 First, we show how to combine financial models with machine learning techniques.
-We used SABR as a complementary information source within our multi-task framework.
+We used SABR as a complementary information source within our multitask framework.
 Second, we developed a hierarchical Bayesian regularization for task embeddings that adaptively determined the level of information transfer.
 Third, our evaluation using Heston ground truth data showed this approach balanced structure and flexibility.
-It provided reliable predictions across market conditions and outperformed standard Gaussian process particularly in data-sparse regions.
-Improvements were notable for longer maturities and extreme strikes, where market data was limited.
-Finally, an application to real SPX market data further illustrated the model’s ability to produce smooth, reliable IVS that respect market observations while benefiting from structural guidance.
+It provided reliable predictions across market conditions and outperformed the standard Gaussian process, particularly in data-sparse regions.
+Improvements were competitive for longer maturities and extreme strikes, where market data was limited.
+Finally, an application to real SPX market data further illustrated the model’s ability to produce a smooth, reliable IVS that respects market observations while benefiting from structural guidance.
 Our work has practical implications for quantitative finance.
-SABR-MTGP provides a more reliable tool for IVS construction, crucial for derivative pricing, hedging, and risk management.
+SABR-MTGP provides a more reliable tool for IVS construction, crucial for risk management.
+Hedging and OTC valuation require market-consistent dynamic models (e.g., local or stochastic local volatility) and Greek validation.
+These topics are outside the scope of this paper.
 By incorporating structural knowledge and empirical observations, it reduced model misspecification risk while maintaining adaptability to market realities.
-Future research could extend the multi-task framework to other structural models or multiple models simultaneously.
+Future research could extend the multitask framework to other structural models or multiple models simultaneously.